by TEI TEI

Admin CHW Pump Replacement

12/26/16 – Contract sent to Tony for approval on this project.
01/26/17 – DK says controls will be done my same team as controls project. Gleason has sent all required paperwork to Tony.
02/01/17 – Project kickoff to be schedule soon. CHW pump size will need to be coordinate with new chiller.
02/16/17 – Kicked off project with Tony, Bert and Gilbert. We are collecting chiller information now and rates for LCC equations. We will choose the chiller during the LCC and design around one chiller during design.
03/09/17 – Pump is on hold pending new size from LCC equations.
03/15/17 – Emailed Dave Gleason for update as work was supposed to begin last week and be done by end of March.
03/21/17 – Tony talked to Dave and they are starting next week with pipe fitting work. Steve Yockey is waiting on TEI to provide new pump size for selection work.
04/11/17 – ThermalTech provides new bypass size and flow for DK to install the bypass.
04/26/17 – TTC checked in with Dave Gleason on progress. Given the cooling season is here, they are going to have to work weekends and off time to fix these issues. The new bypass size was sent to DK but I do not know if it was installed yet.
05/02/17 – TTC checked up on project again with DK for progress.
05/25/17 – Debra stated they will replace the pumps and piping at the same time, later in June.
06/28/17 – Met Debra today on project update and they are getting employees finalized for this project understanding it will be weekend work at this point. Pumps are on order, VFD”s are a new change order, TTC to give to Tony, will be from new project on new PO for up charges.
07/18/17 – Met with Design time to redesign the piping for the new larger pumps on teh CT and CHLR job. Spoke with the manufacturer about sizes and head. Debra submitted new change order for entire project. TTC to estimate the entire job again for Bert with Construction budgets.
08/10/17 – We are waiting for the weather to break to start the pump replacement.
08/26/17 – Debra to drain system and install isolation valves Saturday but this was postponed due to material arrival issues. They will try the week after Labor Day.
09/25/17 – Isolation valves are installed and working. DK needs to insulated still. TTC met with DK and went over remainder of project on site. The scope of the original project lines up perfect with the chiller replacement project except this project does not need to install the bypass piping on the fifth floor. The CHW pumps will be installed by DK in this project, the VFD”s and submittals must be ordered ASAP, and the change orders have been sent for approval to DK and the County.
11/2/17 – Pump replacement work is schedule for December 2017. Debra is currently on site and working on this job.
02/02/18 – Punchlist distributed to Debra Kuempel for project closeout.
04/04/18 – Only insulation remains to be completed. DK committed to wrapping this project up including controls by 4/30/18.
04/25/18 – Debra stated that all work on this project is complete including insulation.  Final paperwork is being completed by DK now for all closeout paperwork.  The final invoices are being delivered to Tony with a cover letter or remaining items.
08/02/18 – TTC inspected all final work including insulation and it looks really good.  No additional punchlist items.  Project is ready to archive.

by GBBN GBBN

CAB – 10th Floor IS & RM Relocation

06/15/18 – GBBN receives phone call to introduce the project and establish the date and time for the Kick-off Meeting.  .
06/21/18 – GBBN meets with Facilities for the Kick-off Meeting.   Project Request Estimate #1826 is to be revisited and updated to current bidding environment.
06/22/18 – GBBN scans all Kick-off meeting handouts and emails to TEI.  A conference call is held with TEI to explain the scope.
06/22/18 – GBBN and TEI discuss scope of new project. TEI submits sub-consultant proposal to GBBN for this work.
06/25/18 – GBBN and TEI work on occupancy number for the training room people for the HVAC ventilation calculations for this work and scope. GBBN provides floor plan for use on this cost estimate.
06/26/18 – Owner sends final occupancy numbers to use for HVAC calculations on this project.  Staff indicated by the owner is CAGIS- 16 and 4 interns, Community Development 4 staff and CLEAR vacant.
06/27/18 – SOSS approved and delivered to GBBN by Facilities.
07/26/18 – TTE sets up backgrounds and prepares job for design for Trades build-out.
10/11/18 – Meeting is held with GBBN, Facilities and Sheriff.  Scope of project is discussed and GBBN requests amounts of existing shelving to be relocated be inventoried by the Sheriff IS Department.
10/12/18 – Amounts of existing shelving, to be relocated, was inventoried by the Sheriff IS Department, and forwarded to GBBN.  Relocation Programming Study from 6/21/18 is to be used.
10/25/18 – Demo is complete.  Abatement is underway.  Carpet tiles are being ordered.
03/28/19 – County has completed abatement, put in new carpet tiles, moved IT racks over to new space. People are moved over to the CAB. It’s an open space for now.
09/26/19 – No project udpates; we are waiting on direction from County on next steps.
01/30/20 – This project has been re-scoped. This version can be archived.

by TTC TTC

CAB – Chiller Rebid Project

03/05/18 – Terry and Colin onsite doing additional fieldwork for redesign. We are investigating new piping connections and existing pipe routes for new configuration. The good news is that some of the piping we thought had to be extended and modified already exist for AHU-1 & 2. The piping changes required on the tenth floor and roof are possible albeit a little tough to do.
03/24/18 – TEI submitted revised SOSS for new design and bid cost for TEI and THP. LCC equation is ready for bidding for air cooled chillers. Revised design is about 50% complete for mechanical changes for the new design.
04/19/18 – ThermalTech release LCC bid documents to manufacturers for rebidding this project for Air Cooled chillers only. Trane, JCI, Carrier and Daikin all invited to bid. The nominal tonnage of the chillers will be 200 tons. Bids and performance numbers are due to ThermalTech April 30 at noon.
04/25/18 – Addendum #1 release today with updated calculations and some other updates to the bidding worksheet.
05/17/18 – The LCC results were published to the bidders, Facilities and Purchasing. Carrier High Efficiency Option #2 was the best LCC chiller for standard refrigerant, model 30XV-2006M40–4-2 using R-134a. And the Next Generation Chiller best LCC went to the York JCI chiller using R-513A model number YVAA0213.
05/23/18 – Terry and Colin perform fieldwork to wrap up new piping changes for new chiller design. We go everything we needed and marked up all the piping on the 10th floor for future changes. We are confident we can do all the piping changes required for convert the building to Air-Cooled Chillers.
05/25/18 – Mechanical drawings are ready for TTC review. Final chiller equipment has been sent to THP for structural design and upon initial review, all seems to be in order. Good news on the piping front, we can find a decent path from the penthouse to the new chillers and route the piping about 41″ above the room and miss the Liebert unit piping rack, found a space to create a walk over the new piping for easy access and can make a new door in the penthouse using existing louver material.
06/26/18 – TTC has marked up final mechanical drawings for editing. Electrical drawings are being updated now for bidding. TTC will get new ITB this week and put the project out for bidding first week of July. I need to coordinate prebid with Bert and John Gilbert.
07/26/18 – Rebid project is on the street. Prebid meeting schedule for August 2 and Bid Opening scheduled for August 23.
08/29/18 – Bid were opened on schedule.  Triton Services was the apparent low bidder on the project.  TTE held a post bid interview with the bidder and sub-contractors on Tuesday at our offices and reviewed the project is detail.  Most everything was covered and Triton appears to understand the project well good sub-contractors for this project.  They would like to get submittals started ASAP on this project.
09/24/18 – Contract appears to be approved and on its way to Triton for signature.  Duke released some data on next year’s rebate program which we should be inline for with this chiller purchase.  THP review the structural proposed questions and returned satisfactory answer to keep the job moving forward.
10/23/18 – PO in for approval and then TEI will schedule the Pre-Const Mtg with Triton.
11/27/18 – Pre Construction meeting was held onsite with Triton Services. Permit response came back and we are working with Triton to get the four responses answered for this job. Terracon must be engaged as a special inspector on the job, and we provided some additional permitting material to take to the city, plus the city will require a General Permit for the structure.
12/31/18 – Two construction meetings so far. Project is underway. Submittals are moving along and being approved. Background checks are almost completed. Triton is mobilizing to site. Crane lift being scheduled for early January. Chiller delivery is February and we are working on getting that moved up. Steel and piping are being fabricated now and that will be the first lift up t the roof.
01/10/19 – Triton directed to get CHLR-3 online first for hot days. Control meeting was conducted with TEI and Emcor on Monday for coordinating control sequences on new chillers. TEI issued updated sequences to Owner, Triton and Emcor for use on this project from this meeting.
01/21/19 – Glenwood found that a 1000A breaker with the correct lug alignment was not available from GE at this time. They submitted a change order for the work. TEI contacted CE power to find a direct replacement for the breaker that will hopefully be less expensive than the submitted change order.
01/22/19 – Control submittals have been received and return immediately. TEI and Emcor worked through specific control questions and TEI issued an updated sketches for 10th floor piping that has to be coordinate with Triton during installation. We also approved a ball valve in the bypass even thought that technically won’t be used until the three way valves are replaced later.
01/31/19 – THP update: Structural steel shop drawings reviewed and returned.
01/29/18 – Const Mtg #4 held. Project is moving along well. Glenwood is setting the transformers today. Piping is connected in 10th floor and along roof to structural steel location. Basement piping is demolished. The 5″ bypass will be started in the mechanical shaft and then routed over the ceilings to the mechanical room. ThermalTech will use Terracon fro the special inspections on the structural platform for this project.
02/17/19 – Steel was erected by 2/17, inspected, fixed and reinspected. Chillers were set on frame on 2/19 without incident. Everything looks good.
02/26/19 – Glenwood pushed to power outage to the backup date of 3/9/19 due to ship date of the circuit breaker being 02/28/19, and Glenwood was not confident that the breaker would be on site in time for the original outage date.  TEI and Glenwood revised drawings and submitted to IBI engineering change for the revised plan.  TEI to schedule a meeting with Triton and Glenwood for pre-outage conference the week of 3/4/19. Const Mtg #6 was held and Triton reported almost all their work is complete. They are waiting on Glenwood and Emcor to finish up the job. the glycol feeder and chemical test company should be on site within the next week for their work.
03/01/19 – Triton reports that glycol feeder is here and being installed. Emcor workers have their background checks in place and will start controls. Glenwood reports after Nester checked in that the breaker should have arrived yesterday and today for sure.
03/11/19 – TEI attended Admin Chiller Startup with County.
03./12/19 – TEI attended Admin Chiller startup for second day with County. Items brought up during startup are weatherguards, heating circuit, convenience outlet, trouble light, local disconnects, additional support on chiller base, additional glycol and differential pressure reader. ThermalTech to start working on these items for resolution. TEI can start the punchlist on the job as well.
03/26/19 – Electrical transformer and breakers installed and turned on.  The electrical final did not pass.  TE and Triton/Glenwood working to resolve issues on project to pass inspections.
03/27/19 – TEI held conference call with Glenwood to discuss change orders. TEI went to site and measure disconnect locations and redoing the work to make it easier for contractor and Owner.
03/28/19 – Barth, Glenwood and Nester met at the building to discuss final location of the new 600A disconnects this morning. Chiller #1 disconnect will be mounted on the West Side of chiller. on Chiller #2 we are basically doing the same thing but we need a junction box and extend the feeders in seal tight.
04/22/19 – TEI releases electrical and mechanical punchlist for final wrap up of project. TEI reviews and gets approved balance reports from vendors and factory for this project.
04/24/19 – Change orders approved for the power items that IBI required.  TEI performs and issues electrical punchlist.  Glenwood on site week of 4/22 making all corrections and CO work with the exception of ground detectors, which are on order but not received to date.  Upon receipt, ground detectors and final IBI will occur. Triton on site doing punchlist items and TEI meets them for new pipe label coordination. TEI asked for Training to be complete ASAP.
04/25/19 – Send Bert change order prices for CT and Chlr. Duke rebate check received and credited to the account, however this money probably can’t be used for additional change order work.
05/27/19 – Glenwood wrapping up change order work and final permitting of electric work for the project. Weather panels are coming very soon. There is a breaker setting we have to coordinate with Glenwood and then we should be able to get our final inspections in the next week or too.
06/27/19 – All construction work in complete. Glenwood set to call for final inspection next week to get final permit closeout for this project. TTC and Emcor meeting on July 9 to do some functional testing of the chillers. Triton approved via change order to remove abandoned and old piping on the project site prior to closeout.
07/25/19 – Triton finalizing all permits. IBI is approved. Triton also performing alternate work for the owner this week. Project ready for closeout.
08/28/19 – Project is complete and final inspections were approved on 8/12. However the owner called with glycol feeder issue with the Pulsafeeder unit. We met last Friday and review the installation with the salesman. They requested we remove all the control interface to the unit and put it back to factory conditions or else they consider the unit “tampered with”. We did everything they asked and then spent another two hours trying to make the unit work with limited success. We contacted the vendor and asked him to bring in the manufacturer to assist. To me this is all warranty work for Triton on this job.
09/20/19 – TTE met Triton on site and verify an AAV on the CHW glycol feed does seem to fix the issue. Triton will install and pipe the overflow back to the glycol tank.
09/24/19 – TTE verified the work is done. This project is ready to Archive.
01/27/20 – Bert called for final closeout confirmation today. TTC checked and all the paperwork and O&M’s are in. However the final payapp is not process. it was found that #10 and #11, balance of contract must be submitted. TTC confirmed with Triton on needed closeout documents.

by GBBN GBBN

CAB – 10th Floor Community Development Renovations

03/26/18 – GBBN contacted by Facilities to plan Kick-off Meeting with Community Development.

03/29/18 – Kick-off Meeting date and time decided.

03/30/18 – Kick-off Meeting held at Room #1022, 10th Floor, of CAB.  Discussions held related to requested scope.

04/03/18 – Facilities sends marked up plan showing clarified scope for SOSS creation.  Photograph of front desk sent later by Facilities to GBBN.

04/06/18 –  Facilities clarifies card reader access system is desired by Community Development.

04/09/18 – TEI sends SOSS input to GBBN for SOSS.

 

 

by THP THP

CAB – Building Envelope Assessment (THP F106)

12/28/17 – THP publish SOSS to Hamilton County as requested.
01/25/18 – Hamilton County approved SOSS.
03/29/18 – Field work to start first week of April.
4/13/18 – THP completed rappelling surveys.
4/23/18 – THP completed survey work from the ground and roof.
4/26/18 – THP compiling data for report.
5/31/18 – THP finalizing draft report for review by the County.
6/26/18 – THP submitted drafts of the Envelope Assessment Report and City Report.
07/26/18 – THP finalized and published Cincinnati façade ordinance report. Signature by Ralph Linne’s required before THP submits report to City of Cincinnati Buildings & Inspections.
08/24/18 THP received letter from City of Cincinnati acknowledging the submitted report.

by TEI TEI

CAB- Auditor #304 BOR Furniture Reconfiguration

Project Lead: GBBN

08/01/17 – Hamilton County requested SOSS to look at Auditor’s request to relocate 4-6 staff.

08/10/17 – GBBN authors and delivers SOSS to Facilities.

10/18/17 – Interior Project Management indicates Auditor has requested some additional work to be performed on next phase.

by TEI TEI

CAB Auditor’s 3rd Floor Furniture Rearrangement

Partner Lead: GBBN

06/29/17 – Facilities contacts GBBN and IPM to request an SOSS for this request via the Auditors Office.

06/30/17 – IPM sends scope and input for SOSS to GBBN.

07/03/17 – Facilities confirms SOSS to include As-Builts and uploading to WebCentral. GBBN issues SOSS to Facilities.

07/17/17 – Facilities delivers approved SOSS to GBBN.

07/19/17 – IPM meets with Auditor Staff to inventory all existing furniture to be relocated and discuss scope further.

07/21/17 – IPM meets with Auditor to present three options for the rearranged layouts. Minor revisions are requested.

07/24/17 – IPM meets with Auditor to present two additional options. IPM reports to GBBN after meeting an additional larger scope of work is to succeed this study and will require an additional SOSS. GBBN reports findings to Facilities to discuss next steps and to clarify titles of each SOSS for clarity in the later phases.

08/01/17 – IPM indicated Auditor has requested additional redesign. Facilities confirmed to GBBN/IPM that Auditor has requested additional scope to project.

08/10/17 – IPM notified GBBN that Auditor now had first rendition of proposed changes for Owner Review. Facilities confirms this information by forwarding drawings to all team members.

10/18/17 – Interior Project Management indicates Auditor has requested some additional work to be performed on next phase.

by TEI TEI

ADMIN ADA Assessment

Partner Lead: GBBN

06/10/11 – Bert Watts indicated this Project was to have an SOSS submitted.

06/14/11 – SOSS submitted in person to Bert Watts.

06/17/11 – This project was not one of the five buildings approved.

by TEI TEI

ADMIN Asbestos Study

Partner Lead: GBBN

10/20/11 – Kick-off Meeting with Pinnacle and Facilities Department

11/04/11 – Facilities Department sent list of buildings to be included in Asbestos Abatement Reporting.

02/01/12 – Pinnacle Engineering requested drawing files for the purpose of documenting findings.

02/02/12 – GBBN sent drawing files to Pinnacle Engineering in PDF form.

02/02/12 – Pinnacle began field investigation at the site.

03/30/12 – Meeting with Tim Schue of Facilities Department to discuss Archibus Input Strategy

04/03/12 – Field work completed, draft report completed and ready for initial review. Discussions ready for drawing production with GBBN.

04/05/12 – Meeting with GBBN and Facilities to discuss drawing format.

04/16/12 – Meeting with Pinnacle, GBBN, Facilities and Abatement contractor to review boiler and gasket removal logistics.

05/04/12 – Meeting with Facilities, Building Manager, Pinnacle and GBBN to kick-off field verification.

05/10/12 – Meeting with GBBN and Pinnacle to discuss drawing format for reports. Tony Matre’s request for ‘General Information Notes to occur on each floor plan are discussed.

10/27/12 – Pinnacle Engineering sent ACM list from Archibus/BRG for purpose of designation conversion from Pinnacle Reports to Archibus entry.

02/26/13 – GBBN inquires with Pinnacle for status of remaining documents to be sent to GBBN.

02/27/13 – Pinnacle reports back that progress has been made and further timetable for receipt will be coming 2/28/13.

06/10/13- Pinnacle Engineering response with email indicating CAB report will arrive 06/11/13 and CAB on 06/14/13.

06/14/13- Galbraith Road Garage Report arrives from Pinnacle. CAB report is not included.

06/26/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

07/24/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/05/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/27/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

02/24/14 – GBBN and Pinnacle Engineering met with Tony Matre to deliver the final reports and electronic copies of the reports. GBBN is completing final report drawings for Pinnacle as final preliminary drawings are delivered by Pinnacle to GBBN. All drawings are to be completed by 2/28/14.

03/26/14 – All asbestos reports and associated drawings are complete. Final drawing input into Archibus/WebCentral is ongoing. GBBN was granted access to the database to complete the drawing downloads.

07/01/14 – GBBN sends another email and follow-up voice mail to Pinnacle Engineering inquiring about status of billings and final completion.

8/29/14- Pinnacle Engineering sends invoices for services to GBBN.

10/29/14 – GBBN sends Tony Matre SOSS invoice status with all of the Pinnacle Engineering invoices included.

10/31/14 – GBBN pursuant to Tony Matre’s request sends letter of documentation regarding scope overruns.

02/20/15 – GBBN in communications with Tony Matre to finalize invoicing on all projects.

by TEI TEI

Administration – Rooftop Computer Room Chiller

09/21/04 – Tony has agreed to put the project on hold and address these issues in the 2005 chiller project upgrade to the 10th floor.

09/20/04 – Building does not have enough power for this equipment. TE scheduled a meeting with Joe Merkt and Tony Matre to resolve.

Design is complete.

Under Quality Control Review

by TEI TEI

Administration – 10th Floor Chiller Upgrade

09/21/05 -This project is being moved to the archive because the whole project is being rebid due to bidding anomalies.

09/15/05 – After much discussion amongst the vendor, the contractor, Tony and Amy in purchasing, it was decided that TE did not use proper bidding procedure on this job and it has to be reibd. TE will provide the new bid documents and prebid and contract time at no cost to the County. TE still has to write an engineering change order for the scope change for the new electrical design (TA to talk to Tony about this new SOSS). The bid and the construction admin will seem to be au gratis from this point forward.

09/07/05 – ThermalTech has received many calls from the vendors on this job mostly in regards to the chiller selection on the project. Trane believes the chiller must be individually test to meet ARI standards and York seems to think that the specifications say that only so many out of 100 need to be test to meet ARI. TTC assured Tony that ThermalTech will review all aspects of the schedule, specifications and bids and provide a written recommendation as to which chiller meets the intent of the specification while maintaining the requirements of the ORC. TTC will ask Tony about getting the purchasing department or prosecutor department to review and provide a decision if needed after the contract is approved. As with any public bid, the bid amount cannot change so the vendors will have to provide the equipment listed to the County at the bid amount or other actions will have to be taken.

09/01/05 – Contracts came back from Prosecutor. TTC to pick up, modify and take back to faculties for approval. Peck and Trane and calling multiple times trying to force us to use the Trane chiller because it is the only to meet the ARI specification. TTC to review ARI 550 and will discuss with everyone after the contracts are signed..

08/24/05 – TTC, TA, AKM and JM talked about how to write the Admin Chiller Contracts and what alternates to take based on our latest discussions with Cinergy and what they are supplying to the building. TTC has the contracts written to accept the EC base bid and Alternate #1 (new service) and for the MC we are taking base bid and Alternate #3 (warranty). TTC stated that the contracts will be dropped off today.

08/08/05 – TE and the County had a meeting today to discuss how to write contracts for this project and what alternates to accept. For the mechanical contractor we will accept base bid and Alternate #3 for 5 year parts and labor warranty. In discussing the scope of the electrical alternates, it was discovered that to accept the new feeder to the 10th floor does not truly provide the full extent of the County desire to have backup power from both grids to building at all times. It is a convoluted situation but basically the building now has 8000 amps to the building from 2 separate grid feeds. If either grid goes down, the building can still maintain power from the other grid. The max usage right now is 3700 on 95 degree OA temperatures. If we only feed the new chiller with the new service, the main building will have no additional power savings. In the end it was decided that the best thing to do was to try to have the contractor provide a new 1600 amp service into the basement from both grids and provide a transfer switch in the basement as well. This will slightly modify the original scope of the Alternate but will handle TE’s misunderstanding of what the County requested in the design process of this job and in no way would have changed the outcome of the bid process. TE will research the changes and advise the County on possible cost and acceptance by Cinergy.

08/01/05 – The bids opened last Thursday. TE is working on which recommendations to accept for this project. The numbers varied for using bus duct in lieu of standard conduit and feeder from a $400 deduct to a $127,000 adder. Because of this TE has to do some research to see what is correct. TE believes that we will be able to accept Alternate #1, Alternate #3 and Alternate #4 and still be under budget for this job. We want to write the contract on either Tuesday or Wednesday of this week.

07/20/05 – Tim Centers from Habegger called to ask about freeze protection on the chiller since we aren’t using glycol at the Admin building. TZ is on vacation but we are using heat tape and DDC alarms to protect the chiller and piping. Ryan Halverson from Trane called to ask about removing the word “Scroll” from the schedule so everyone is bidding “Screw” compressor chillers and TTC will do that in the Addendum.

07/18/05 – TE to work on Addendum 01. Fix the bond issues for purchasing. Move rigging time to full weekend. Move conduit pull to weekend. Add01 should go out on Tuesday.

07/12/05 – The bids are out and the prebid is set for 07/15/05 at 1:30 p.m.

06/15/05 – JM reported that discreet chiller control will be fine for this job.

06/09/05 – TE emailed Tony and Joe about using discreet points to run the chiller plant instead of adding a Trace interface. TE needs to wait for approval from Tony

05/20/05 – TTC dropped off the chiller plans for review. The specs are not complete yet for review and TTC has not decided the best way to bid the chiller (LCC, Alternates or Base and Equals). GD has returned his comments to TTC.

04/27/05 – TTC, TA, TZ, Joe Merkt and AKM met to discuss the project so that TE can design in a way that doesn’t go too far over budget. The cost estimate turned in last year only had $33,000 for electrical upgrade. The plan is to make or use the mechanical scope adjustable so that some additional monies can be diverted to upgrade the electrical. TE has to do more fieldwork to check on the piping arrangement so that TE can fully comprehend the additional return pipe that goes from the 10th floor to the basement. Basically the 3-way valves have a return system that goes from floor to floor UP the building and from the 10th floor back down to the basement.

04/22/05 – This week THP has delivered final structural dwgs to TE for project. TE finalized a control strategy within TE for the new chiller and discussed hydraulic issues of existing piping and pumps. GD gave TZ and TTC direction on sizing the new 10th floor pump. TTC decided that JCI controls can be modified during the upcoming project and that existing control work on 10th floor chiller would have no adverse effects on new system. Chiller S/S and Pump S/S and statuses can be re-used and extended as needed during 10th floor project. Trane sent project info to TE on the existing chillers: Evaporator, 45-55 F, 480 gpm, 2-pass, 26.5 ft wpd, .0005 fouling factor, Condenser: 85-95 F, 581 gpm, 2-pass, 21.7 ft wpd, .0005 fouling factor, CAB-CH-1 TRANE CENTRAVAC CVHE-020B L85B26377, CAB-CH-2 TRANE CENTRAVAC CVHE-020B L85B26376.

04/05/05 – TE coordinated with THP on steel locations and scope for new air cooled chiller. THP will meet onsite with Joe Merkt to field check existing steel conditions this week. THP requested $2000 for fee’s for this project. TE approved. This cost wasn’t in the original scope of fee’s but TE will try to incorporate into project but will save for engineering change order if needed.

03/21/05 – TE kicked off project with a site visit for TA, TTC and TZ. SRB visited the site for equipment and piping takeoffs during the same morning. TE began talking to vendors about equipment weights and chiller selections. TE discussed the cost estimate and various upgrade options at the building for this project.

02/09/05 – Received Signed SOSS

01/12/05 – Project Kickoff meeting with County. TE to begin design.

Project Status

03/24/05 – TTC sent the revised change order to ICS for him to sign and prepared it for mailing, stopped them and reprinted them again because Gary called Nancy to schedule the City and Fire Marshal inspection to find that cost is $475 additional to the city fire inspection so now I am advocating that the change order be re-written back up to $1400 to cover the additional charge and Gary will just absorb the rest of the overage from his profits. Larry Keifer has required a FULL BUILDING HORN & STROBE test. Joe called TTC to express his disbelief that a full building test would be required and indicated how difficult that would be to do in one hour. TTC sent an email to Larry Keifer requesting clarification and hopefully explanation on this upcoming test. Tentatively the test is scheduled for Tuesday March 29 at 6:00 a.m.
03/23/05 – ICS finished relocating the fire alarm strobe for our new fire alarm permit. Joe tested the strobe and it works. TTC talked to ICS about painting the wall where the old strobe was moved and negotiated a $450 reduction in the original change order down to $950 and in return the County can touch up the column at their convenience. TTC and JM agreed that this seemed reasonable. TE will rewrite the Change Order and get ICS to sign it for final billing. ICS has to coordinate a final inspection between Larry Keifer and Nancy at the Cincinnati Fire Dept. He hopes it will happen this week, but TE believes this will take at least a week to get scheduled. ICS is to coordinate with Joe Merkt when FA test is actually scheduled.

03/15/05 – ICS submitted a change order price of $1400 to move the strobes for the approved fire alarm permit. Diana approved. TE will now scheduled (we have already called to schedule the work) and assist in the installation of strobes as needed.

03/08/05 – City Permit Office reported today that Fire has approved and they are in the process of setting up the permit. They will call me at TE when ready.

03/07/05 – TTC called the City Permit desk today. Plan Examiner approved on 02/22/05. The Fire Department has the permit now for approval. TTC called 352-4314 and convinced them that the Fire Sprinkler work was completed under Permit # 2004P06759. The Fire Inspector updated the database. Now the city should call us in a couple of days with the final permit price and tell me it is ready to pick up. TTC will call ICS to get them to pay for the work and give us a CO price to do the work for this job.

03/03/05 – TTC contacted the city permit department about status – no word yet.

02/22/05 – TTC contacted the city about permit status – no word received back at this time. We will continue to push for final inspections. TTC authorized ICS to bill for $3500 leaving a $200 balance on the project and whatever this change order turns out to be since all of the original work is complete and punched.

02/14/05 – TTC requested final paperwork from Carrigan again. TTC requested allowance usages from B&J on project for final bookkeeping.

02/11/05 – TTC contacted Preferred about lowering the sprinkler head in the public work area. ICS called to report that the final IBI inspection was today and that everything passed. Final paperwork is en route. Preferred called back and will get it taken care of quickly through Joe Merkt.

02/11/05 – TE submitted the Fire Alarm Permit in person today. The FA permit number is 2005P01862. It will have to be paid for when it is picked up. The review should take under 12 days. TTC asked for Bill Spurling to do the review. TE will keep everyone apprised of progress.

02/10/05 – TE has finished recreated 4 sets of new fire alarm permits for the microfilm area. MS talked to the building department today and we can try for a mini permit but the drawings have to be on 8.5×11’s only. These plans are on 11×17 and stamped and signed so we are going to submit them as a standard FA set and not walk them through as that doubles the cost. ICS will have to move one strobe still to meet the new plans we are submitting.

01/31/05 – TTC talked to Bill Spurling at City plans exam. He did not know where Larry wanted the additional strobe – perhaps the darkroom. He asked me to talk to Larry, so I placed a call to Mr. Keifer and when he calls me back, I will get to the bottom of what we need to do and get it taken care of immediately.

01/27/05 – TE mailed Gary @ Integrated blank final document forms: G706, G706A, G707, Affidavit of Contractor Prevailing Wages and, Affidavit of Original or Sub-Contractor.

01/25/05 – City is now requiring a possible new strobe in the public area which would mean we have to take out a fire alarm permit. Stiles still believes that you can get a TCO for the space and we can pull that permit and get that work done and still make the city happy. TE has placed two calls into William Spurling, the city Plans Examiner, to discuss the issue with him. TE will keep the County apprised.

01/21/05 – Project passed above ceiling inspection with IBI. EC is 95% complete. Larry Keifer may still require a Fire Alarm permit to be pulled for this job although TE doesn’t believe that is necessary. TE coordinate with Preferred to add a sprinkler head in the vestibule as it was shown on the drawings but not to add the one above the microfiche charts as it would tend to get damaged from the anticipated use plus there is not 18″ clearance above the racks.

01/18/05 – Plumbing is finished including vent. HVAC has installed toggle switch for exhaust fan. Auxiliary exhaust fan is burned out so TE instruct B&J to buy a new motor and replace it. B&J has the motor in stock now and will install it on Wed Jan 19th.

01/12/05 – Dennis Grimm of Carrigan & Grimm contacted, requesting 3 additional documents (Affidavit of Original or Subcontractor, Affidavit of Contractor Compliance-Prevailing Wages and Consent of Surety to Final Payment) to complete submission of all final documents. Dennis not sure how to obtain the forms therefore the forms were mailed to Carrigan & Grimm on this date.

12/27/04 – TE contacted ICS about catching up with GC. They said they would be on site this evening at 4:00pm to catch up. ICS will also begin installed lead anchor bolts with 1/4″ #20 eyelets for Communications to hang their J-Hooks for data cable support from. TE advised them not to exceed $640 dollars for this work. Gary Kercher agreed to this pricing scheme. DES can contact Mitch that they will be able to move their cables to the new supports soon as this will need to be done before final inspection.

12/10/04 – TE advised the GBBN can create a wooden soffit around conduit and panel box in vestibule and paint the wood and the cover of the panel and still meet code requirements for the panel. County arranged to have carpenters do the work through a work order. Matthew shall enter a work order into the system.

12/07/04 – Construction Mtg addressed the scheduled as we are about 1 week behind due to holidays and with Christmas approaching, the project is likely to stay a week behind. County asked TE to look at cleaning up existing panel in vestibule for better aesthetics. PC is complete. HVAC will be in on Thursday night to relocated controller for hood fan.

12/01/04 – GBBN & TE resolved the RFI requested by ICS electric in regards to the existing electric panel in Room #201 near the countertop. Microfilm has agreed to redesign the countertop to accommodate the electric panel. GBBN will follow through with 3R and TE will notify ICS that no additional work will be required on their part.

11/24/04 – Rough-in inspection complete. Inspector may request clean up of existing cabling above ceiling. TE to investigate.

11/03/04 – Asbestos Memo was issued to contractors for tile removal by pinnacle communication. PDF. Preferred (Jeff) is confirmed for evening work tonight (11/03/04) and we have coordinated with Joe Merkt for this evening work. TE has contacted all contractors about not working on 11-04-04.

11/01/04 – Carrigan called today to discuss existing piping layout, mixing valves, and hot water heater placement. It appears that some ball valves were added to self regulate temperature at the film equipment. After discussing the situation with Dennis, we believe it would be better to pipe the building hot water into the water heater and have it “boost” that water up to inlet hot water temperature at the mixing valves thereby giving the process room immediate hot water when needed. In it current configuration it backfeeds the building hot water system which would give us unpredictable results. TE will confirm this at tomorrows meeting but in the meantime, we have asked Carrigan to give us a price to replace both temperature mixing valves with new Simmons Temptrol mixing valves that have integral temperature sensors for visual tracking outlet temperatures and piping the water heater in series with the buildings hot water supply to the process area. We will discuss all this at tomorrow’s meeting.

10/27/04 – Fire Permit has been approved: 2004P06759. Carrigan & Grimm schedule of values received and approved.

10/26/04 – Code questions were received by permit office. ThermalTech responded with these answers. PDF

10/25/04 – Carrigan asked to rework the vent arrangement on the sink and equipment in the new area to prevent a flat vent condition on the floor drain. TE has approved this routing of the new vent. Minor change and not cost to the project.

10/19/04 – Const Mtg 2 held. Plumbers, HVAC, Electricians, Fire Protection on board and ready to begin evening work. Scope coordinated through GBBN during project meeting. Owner agreed to allow ceilings to be removed during remodel work and replaced at very end as long as contractors take out the minimum required for each phase.

10/07/04 – Const Mtg held. Carrigan & Grimm was not present. All work appears ready to start on October 18. Phasing and project schedule mostly approved. Preferred needs security checks approved and DES will send them to JJ. We may see additional monies requested from EC for additional permit cost. We may or may not approve these additional funds depending on how the project goes. We can use allowance money if needed and nothing else comes up. TE will post minutes after they are received from GBBN.

by TEI TEI

Administration – DDC Upgrade to Metasys Gen III

02/27/06 – JCI reported today that all final documents are on their way to Tony and they will train the Admin staff on outside access for this project and take care of small punchlist items.

02/23/06 – Final paperwork received. No significant punchlist items other than to answer a couple of questions for us on the control system parameters on the job. TTC will closeout the project. TEI UPS final payapp, closeout paperwork and O&M’s to Tony Matre today. Final Punchlist items included: Item #1: What controls the EF-1 operation? Is it just the rooftop units? That appears to be the only thing telling the fan when to run. The problem is that aside from the OA enable setpoint that EF status should be tied to all the air handlers for relief during economizer mode not just the RTU’s they are almost incidental to its operation. Please clarify. Item #2: The speed drive has been hunting continuously since Mike Asbrock made the latest programming changes. Please investigate and remedy. Item #3: Are you providing a laptop on this project for Joe Merkt? Regardless, we need training to show Joe how to pull the ADS server up from outside the office. Have you coordinated with HamCo IT for internet access or port security control into the building? Does Joe and his staff understand how to view from a remote location?

01/18/06 – Wayne Jackson reported the JCI has been on site working on the punchlist this week. Joe Merkt confirmed they had indeed been at the admin building. The licensed admin software is not on site but JCI will find it and deliver it to Joe. A full backup of the JCI system was completed today. TTC will schedule final punchlist visit for next week to wrap up the job. JCI will provide final documents and billing by end of January.

01/10/06 – TTC emailed JCI to inquire about taking care of punchlist emailed on 12/13/05.

12/13/05 – TTC emailed a partial punchlist to JCI. No response has been received or questions answered yet. AH-6 has been repaired by JCI at this time.

12/06/05 – TTC stopped by the building and began the punchlist today. It is not complete and a second trip will be needed. TTC to schedule in the near future.

12/02/05 – Admin Punchlist has been scheduled for 12/06/05. TTC will distribute soon thereafter.

10/24/05 – JCI is complete and TTC is supposed to contact Joe and setup final punchlist for the project.

10/04/05 – JCI was supposed to complete the job on Friday. TTC will check and find out and will advise the County.

10/03/05 – JCI did not get finished with the DDC project as promised. They believe they will be done this Friday with graphics and 10th floor upgrades 10/07/05. TTC has to follow up with the punchlist as soon as they are complete.

09/19/05 – JCI emailed back and confirmed that the deadline is 10/01/05 and they will be done. TTC told Tony that TE will punchlist the job and get punchlist complete by mid-October.

09/16/05 – TTC emailed JCI about working onsite and meeting the 10/01 deadline.

09/08/05 – JCI emailed to commit to an 10/01/05 completion date on all remaining items on the DDC project. Parts should be in late this week and they can start this week or next. They will contact Joe for on-site work.

09/01/05 – TTC emailed JCI today about final completion again.

08/31/05 – Joe called about getting a final date for completion. This project has gone on too long and Tony wants a completion date that includes the change order cost.

08/24/05 – TTC contacted JCI with all the change order and allowance usage information to get final approval for doing that work. Once complete no allowance money will be left and JCI should bill Joe directly for remaining cost. TE will advise and information comes back.

08/23/05 – JCI and Wayne Jackson have sent in a revised current sensor change order without all low voltage wiring ran in control. The new cost is only $5700. TTC passed along to County for approval. TTC had a voicemail from Mike Asbrock on Monday stating that the hot water pumps were alarmed but not the boilers per say and the only way to know the heating boiler was in alarm was by a decrease in hot water temperature.

08/08/05 – JCI called in response to my inquiry about project status and provided the following report. The project is still about one month away from being complete. Each floor has a reset schedule that they were unaware of that appears to have been added after the original programming and they have to go back and put that into each DX unit. The will have this done in a couple of weeks – their technicians is on vacation this week. The graphics aren’t started yet but they will have their technician do that in a couple of days – but they have to completely wipe the 630 main program off and start over to make it right. They are working on re-pricing the current sensors because their electricians are 100% complete at the building and to bring them back that will cost mobilization, but I told them we still need a price and to coordinate with Joe the scope of what he wanted on the closed feedback controls. Wayne Jackson went on to say that they know they still have a lot of small cleanup work to complete and I told them that and to update the as-builts with all the new sequences.

08/03/05 – Joe called today and asked about revisiting the circuit sensor change order since the chiller job came in under budget. TTC sent JCI and email about final status and about what the cost would be to upgrade the controls and sensors to monitor actual running status of the equipment at the Admin. The old control system had open feedback also – so it didn’t get upgraded during the project. TTC also inquired about the graphics upgrade and when the project would be complete.

06/28/05 – TE received invoice # 01 from JCI but as Pat pointed out, it is not on an AIA form and it doesn’t include the correct paperwork for approval. TTC will contact JCI.

06/15/05 – Joe called today to report that Tony will not be accepting the current sensor change order. JM also reported that JCI was finished with floor 1-8 on all HVAC units. No graphics have been installed to date. TTC said that made sense because that will be the last part that Ed Braemer puts it when he arrives on site for commissioning and downloading.

06/09/05 – TE received update submittals showing the chilled water system for approval.

05/19/05 – JCI is onsite with ACPI on the 8th floor working on the FCU panel. A new DX9100 is being installed and ancillary equipment is being installed before the changeover takes place.

04/22/05 – TE has reviewed submittals and can approve after JCI answers a couple of small questions about AHU-1 and 3. Should leave TE on Monday.

04/06/05 – TE received submittals on controls from JCI. TE will begin review of entire packet for use on the new job. TE is still awaiting direction from the County on whether to accept the CO request for fan status on each air handler.

03/24/05 – Submittals to TE at end of next week. Electrician (ACPI) has pulled the N2 bus around the building. JCI thinks that they will be on site a couple of weeks after that. JCI is still awaiting an answer on circuit sensors.

03/03/05 – JCI turned in a $$9226 Change Order request to add current sensors to all the air handler fans. Apparently the existing system only has open loop control of these fans meaning that if the fans are commanded on, the system shows them as running, if they fail or don’t turn on, the DDC will not show them in alarm. Please advise on how to proceed on this change order request.

01/19/05 – TTC, Ed Braemer and Joe Merkt discussed additional air handler today but couldn’t resolve the numbering issue. We will meet onsite on Jan 24 to resolve. Joe believes that the quote includes all air handlers for this project.

01/12/05 – Joe Shutte called to tell us that an additional air handler on the first floor was discovered during field work that does not appear on any JCI control dwgs or as-builts, therefore it was not included in the state term price given to us for this project. TE informed them to match the existing sequence and give us a price to upgrade it on this job pending a decision from Facilities. We are awaiting the allowance/change order price.

11/05/04 – Pre Construction Meeting

09/20/04 – Tony and Diana requested an SOSS for this project today. TE completed the state term contract for JCI and has already sent it down.

by TEI TEI

Administration – 6th Floor Sink Addition

03/16/06 – Mark reports the sink is compete. I am set to archive this project at the end of the week.

12/28/05 – TTC called Blue Chip to find out the status of the permit application.

12/14/05 – Permit drawings have been signed by ThermalTech and sent to Blue Chip for submissions. Copies will be delivered to the County on 12/15.

12/09/05 – TTC contacted Blue Chip about acquiring this permit for the County. Sam completed a permit drawing for use on this project. TE will get 3 stamped sets of the drawings for Blue Chip if they agree to get the permit.

12/06/05 – THIS PERMIT WAS ACTUALLY FOR THE ADA TOILET UPGRADE. Wayne called to say the scope changed on the project and the only thing they need is a permit for the sink, TTC reported that the permit was pulled by Blue Chip Plumbing already. TTC called Blue Chip and is getting a copy of the permit faxed over to TE. I will scan it and fax it down to Wayne when we receive it.

11/07/05 – The County has decided to the do the work in house and TE will assist in getting a permit for this job and reimburse the cost through Blue Chip for the permit.

10/03/05 – TE sent out a project update but we have not received a signed SOSS for this project yet.

09/23/05 – TTC & MS are meeting the County on site to discuss the scope of the project. We will then write the SOSS for the project.

by TEI TEI

Administration – Engineer’s 7th Floor Renovation

12/26/06 – This project has been put on hold for so long, it is being archived by TTC.

12/14/05 – TE finalized cost estimates for the new phase I work per Cole and Russell’s revised scope. Archibus is up to date. The HVAC scope remains the same because the areas affected are all within our HVAC scope. The electrical is different and Mike Stiles will update the Phase I cost to reflect the smaller scope.

08/05/05 – Diana sent an email stopping all work on this project until the Engineer’s Office releases everyone to do more work. ThermalTech is nearly done with the design at this stoppage point but will hold all work until further notice.

by TEI TEI

ADMINISTRATION – AUDITORS COMPUTER ROOM & RENOVATION

12/26/06 – This project is being moved to the archive until such time as the tenant comes up with the funds to restart the project.

10/01/06 – This project has been put on HOLD till further notice.

09/01/06 – Based on our new chiller design, the expensive computer room unit required for this room could easily be handled by the new/existing 24/7 chilled water system in the building. This may save the Auditor $10,000 whereas the only thing they would buy now is a fan coil unit and the piping from the CHW system to the space. They would be our first tenant to jump on the new system.

07/25/06 – After the County decided to pre-purchase the equipment, TTC has been working with the Liebert rep to decide which equipment would best suit this application so it can be purchased. TEI has not received data from the Auditors office for loads for the new equipment yet.

06/09/06 – TTC sent contracts down for approval on 06/28. Mark Donnelly called today to say that the Auditor would like to pre-purchase the Liebert unit and UPS for the space. TEI will put together a small project pre-purchase spec and bid this out for delivery in November.

06/14/06 – TE and CRA had a meeting to discuss cost estimates with Auditor and even though a miscommunication of what they wanted estimated prevented the meeting from being a total success, there was enough good information to allow us to redo the cost estimates by 06/21/06.

05/29/06 – TTC emailed Liebert about equipment options and budget cost for the upcoming cost meeting. Our fieldwork at the last meeting located the condensing unit on the small roof adjacent to the atrium roof. The architect will have to cut an access door into the wall in the storage room where an old window use to be. Spencer, TTC and Joe Merkt discussed options on this door. TTC stated that when this new condensing unit was installed that the GC will have to install a set of handrails on the outside parapet since the unit will be within 10 feet of the outside roof. Joe seemed okay with the location and this will give us 24/7 cooling without using city water.

05/25/06- Kickoff meeting left TEI with two cost estimates to assemble. #1 leave the computer room on the fourth floor and estimate to the reworking of personnel on 5. #2 move everyone up to the fifth floor including the fire protection system and servers and UPS. We will meet in two weeks to go over cost and issues for design.

05/22/06 – Mark setup a kickoff meeting for this project on 05/25 at 10 a.m. SOSS’s have been signed.

04/07/06 – TTC and GM attended a meeting with the Auditors (Paula Drake, Terry Munz, Stu) and the County (JM, Jim Eby, DES) to discuss the future of this project. DES stated that the job will most likely be incorporated into a 2006 pop up project to move the computer room up to the 5th floor. The Auditor has an offsite computer farm that consists of 6 servers that they are paying $4000 a month. These may be moved into the basement to save money. We will try to design a new project asap to get a 5th floor computer room created. We will reconvene on Monday with Spencer to discuss options.

03/13/06 – Mike Stiles emailed the County for clarification on the scope of this project:

I would like to confirm the items expected under the SOSS and address what it fixes and does not fix. As you know from previous conversations there are several problem with the computer room as it pertains to power. The SOSS written and approved covers some of them but not all, as discussed. The first and most immediate problem is the power strips on the back of the racks failing because of the poor design of the device and loading. The SOSS is to replace the existing power strips, add new power strips on all the racks, and double the amount of circuitry. This will allow for better loading and add redundancy for dual power supply devices.
This will fix the immediate problem but I would like to clarify that this doesn’t to fix the next main problem. The current UPS is loaded to full capacity (82%-95% which is greater the manufacturer recommends). The additional circuitry will fix the loading issue at the power strip but not at the source. No additional devices should be added to the computer room currently even after the additional circuitry, without the UPS being upgraded. There is an option but not a great one. Some of the new circuits could come from building power, not the UPS, and will allow for more devices being added. The down fall of this is the devices connect to these circuits will not have conditioned back-up power. I need to know if the County would like all the circuits to be from UPS power or a mix of building power and UPS power.

by TEI TEI

Administration Electrical Device Identification

02/02/07 – DK has submitted all final paperwork on this job including the final pay application. TTC believes it is okay.

01/17/07 – Debra-Kuempel called the County because they are not turning in correct payapps. TEI will called and send over the final correct pay application for them to turn into TEI for approval. TEI has faxed a final version to DK for use. Alberta received it.

01/10/07 – Project completed in 2006

10/30/06 – Request to cancel change order was accepted. NR returned pay application #5 because of discrepancies in pay amounts. Final pay request should be for the remaining $4,711.14 with no change order.

09/11/06 – TTC contacted Steve Kallmeyer about change order cost on this project. We are trying to get all change order within the $5000 allowance.

07/06/06 – Close-out meeting held on location. DeBra-Kuempel still needs to put the new panel schedules in the panels; the old schedules are to remain in the panels. There were no other open issues. DeBra-Kuempel provided a CD with the panel schedules on them. I will have the drawings together within the next 4 – 6 weeks. TE will have (3) full sized sets delivered to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Joe Merkt). TE will make (3) binders of the panel schedules and deliver them to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Joe Merkt).

06/21/06 – I spoke with Steve Kallmeyer about devices that were not identified. He is sending an electrician out to the jobsite on Friday, June 23, to finish the identification. He is to call me on the following Tuesday to set up a close-out meeting.

05/09/06 – DK sent application #2 with wrong numbers in contract. Job is being closed out this month.

01/09/06 – DK began working on device study in building.

11/15/05 – Contracts have been submitted.

by TEI TEI

Administration – ADA Lighting Upgrade

03/17/06 – TTC walked the site today and saw two lights on the 6th floor being installed. One worked and one didn’t. There appears to be a problem with some of the light fixtures. They may have to be returned for work at the factory.

11/03/05 – TTC sent out the final Addendum for the new bid on this project. TTC took out an ITB since the cost estimate was above $25,000 on this project. The bid is set to open on November 10th. I don’t know why I was rushing to send out addendum today – I actually had till Monday. I reset the estimate to $62,500 after talking to Mark and included a lamp for each fixture as part of the bid. Add01 – PDF

07/20/05 – TA talked to Mark about MS working on the ADA project. MS will provide a new update soon.

02/22/05 – TE is waiting on further direction from the County on any additional work required at the site. Phil and Christine are creating a detail for the additional lights in the lobby area on floors two through six for future use.

02/04/05 – PL visited the site with Brian Meyer and MWH and readjusted some of the lighting in the lobby. PL dropped off the refurbished fixture for the CH too. Rebecca Groppe did not like style and installation of the lights, but TE thinks it looks great. We believe that the space appears open and clean. If there is anything that TE can do to make things better, please let us know.

Phil Lambing has been onsite surveying existing fixtures and will devise a plan to upgrade the lighting appropriately.

by TEI TEI

Administration – ADA Toilet Upgrade

03/16/06 – Mark reported that they are waiting for partitions to come in right now for installation.

11/02/05 – Blue Chip called to check status of permit. TE stated the invoice could be sent directly to Mark Hardin. TTC does not know current status of doing this work inside the building by trades, but will check with Mark to find out.

09/07/05 – Project is complete. TE has been instructed to assist in the plumbing permit retrieval. TTC contacted Blue Chip who will turn in for the permit for the County. County trades will do all the work. TTC will print out 3 sets of 11×17 drawings provided by Ben at GBBN and give them to Don Grace for city application. Blue Chip will be the county for their time.

by TEI TEI

ADMIN – RECORDERS OFFICE CUBICLE RENOVATION

03/20/07 – As-Built are here. TE will return this Thursday to MD. Closeout is complete.

03/15/07 – Permits are closing out and job is complete. Ready for Archiving.

12/26/06 – TTC has given all dwgs to contractors and owner for work on this project. The FPC does not need a permit for the limited work and that is to be scheduled directly with Joe Merkt. The IBI was acquired for this job and faxed to Mark Donnelly to give to Jim Eby when need for final inspection.

11/28/06 – TEI was originally working for CRA on this project but the scope got big enough that we took out our own SOSS. Today we reported the J2 fire protection submitted a quote 2 weeks ago for $400 to do this work and that no permit is required to do this work in the space. It is our understanding that J2 can be contacted and coordinated by MD and JM to get this work done in conjunction with County Trades modifying the space. TE submitted for an IBI for this project and scanned the permit and sent it to the County for use on the electrical side of this project.

by TEI TEI

ADMIN – 10TH FLOOR CHILLER REBID

05/03/07 – After we found that the existing expansion tank was ruptured we did some super quick engineering and worked it out with PHB to install a new bladder tank Expansion tank on the 10th floor set to 20 psig. After some initial heartache, we found this fix to be working out nicely.

03/15/07 – Joe reported that pump head was extremely high in low load conditions. TE to investigate.

02/19/07 – PHB has finished the pipe bridge work and corrected the reverse return system at the building. First test appear to have everything working correctly. TEI directed everyone to closeout the project.

02/02/07 – PHB is having Bob Gettler’s brother come down to the building next monday to look over remaining work to take Bob’s place.

01/17/07 – TE distributed the punchlist and informed contractors to turn in all closeout paperwork and final pay applications for this project. Punchlist PDF

01/05/07 – TE has to get Trane and PHB to fix a couple of items on the chiller installation and some prices to finish the work including a valve in the pipe bridge and some steps.

12/04/06 – TTC informed contractors to use 10/11/2006 for warranty start date on this project. It was the weekend of the startup and electrical shutdown and will be as good as any to use as the chiller was brought online for service shortly after that. After talking to Joe, we may want PHB to give us some prices for steps and rails up to the chiller control panel as well as the other RTU’s with our allowance money.

12/01/06 – TTC talked to QCE and they will have O&M’s here in a week. TTC has approved payapp #2 for this project for them and sent to Tony.

10/25/06 – PHB Change Order for piping change will come out of allowance of $5000 which has not been used yet (see below). Queen City also confirmed that the $5000 allowance was not used to date.

10/10/06 – TTC check on weekend shutdown and everything seemed to go well. The chiller was not actually started. Today, Bernie told TEI that he believed QCE would have a person down at the Admin building to actual provide power to the chiller. This will allow the chiller to be tested and factory started on Wednesday as planned. TTC may stop down in the morning to check on progress.

10/02/06 – Per everyone’s agreement the chiller will be started on 10/11/06.

09/01/06 – Greg reported today that Duke has been inside the vault and working on the new service. This could mean a completion time of next week. QCE will follow that up with a scheduled shutdown of the building (after talking to Joe) for their tie-in. PHB will be down next week to finalize the piping bridge crossover removal work and then we can really wrap this project up. The new chiller will need to be powered up and tested and the controls (current status unknown by TTC) will have to be checked. Any new computer room system can be added to the system from that point on as long as everyone understands that chilled water will now be available in the building 24/7 either my the main plant or the rooftop plant.

08/18/06 – TTC emailed Joe and Tony costs on converting the system to 24/7 chilled water for computer room units. It can be done with little up front cost to the building and each user would pay for connection to the system. Each computer room unit would get a tap on its associated floor, a check valve, an isolation valve, and an inline pump to provide the chilled water to the unit.

06/29/06 – TTC met Bob Gettler on site to discuss the removal of the bridge connecting the supply and return piping for the CHWS/R. PHB will put a price and time together to do this as Joe and I decided it should be done on the weekend and some equipment should be moved to keep it from getting wet. They will put a price together and send it to us for approval.

06/08/06 -TTC brought all the equipment to the site for field work. We took many readings, some of which don’t make sense. The pressure drop across the flow meter in the basement is showing a chiller gpm flow that is 100 gpm higher than design with or without the auxiliary pump on. The pressure drop at the top of the riser is about 30 psig but the pressure drop across the coil is 1.5 psig. The water temperature at the inlet to the coil is only 57.4 degrees. Cooling is an issue on this floor. The bridge could have something to do with it. TTC requested data from Trane and the balance report to help in resolving the new temperature issues. TZ remembered that during design, this floor had temperature issues during hot weather too so that may be indicative of a long term problem.

06/06/06 – TTC and Joe have been working on AHU-10 serving the 8th floor space since the flow and temperature don’t seem to be exact. Joe also showed TTC a piping bridge that existing on the existing supply and return piping that wasn’t shown on the TEI PID. After consultation with Todd Zurweller and Gary Davies, it was agreed that the new design cannot work without that bridge being removed and blocked. TTC will stop by on 6-7-06 to field verify flows and pressures. TTC emailed Tony with Duke Energy contact for new electric meter application. QCE is negotiating with PHB on roof curb space for new conduits. TTC emailed PHB about removing the 10th floor bridge.

05/26/06 – The steel beam has been moved to its correct location. TTC gave PHB the okay to remove the expansion tank as part of the allowance as long as it is under $1000 as we have a $5000 allowance on this project for the MEP contractor. This expansion tank will no longer be needed on this project but it was not shown as demolished on the plans – it was left as is. TTC contact QCE to have the remove two conduits that are in the way of finishing the demo of the existing chiller.

0523/06 – The chiller lift went very well aside from the new support steel being 1-1/2″ off from where it should have been welded. THP, PHB, TEI and HamCo met onsite to discuss options. TTC instructed PHB to move the steel to the correct location. THP provided a field sketch showing how to use a spacer to keep extra work down to a minimum. PDF. The old chillers were removed and the new chiller set in place with little other incident. PHB pointed out about the expansion tank in the 10th floor. It is not necessary, Joe would like it removed, but the TE dwgs do not show it being demolished. TTC asked PHB for an additional cost for this work. PHB pointed out two conduits they will need removed and TTC said he would get that taken care of by QCE. The electric penetrations for this job were also discussed and TTC said that QCE could use the refrigerant penetrations to save additional roof penetrations and TTC will talk to QCE about this. Lastly, PHB asked about routing the new chilled water lines out of the mechanical room in a different spot closer to the steel. TTC approved this new location as he did not see negatives to the proposed locations.

05/18/06 – PHB and TEI are discussing Saturday Chiller lift. PHB asked about vibration isolators. They do not come with the chiller but TTC would like them to be put under the chiller before it is set on the ground. We asked for a cost. It should only be a couple of hundred bucks.

05/09/06 – Const Mtg 2 was held and mtg notes distributed. The chiller list is scheduled for May 20.

03/29/06 – TTC distributed meeting minutes from yesterdays meeting. We are shooting for do the weekend shutdown over the May 20 weekend if the chiller arrives the week of May 15 like it is supposed too.

03/20/06 – City asked for stamped structural dwgs on this project. TTC had THP stamp their drawings and send it over for PHB to replace in the permit set at the city.

03/15/06 – TEI approved chiller, pump, wiring, mat and switchgear submittals and sent them out to the contractors and owner3/09/06 – Chiller and Electrical submittals have arrived. The chiller selection will be modified for 25% glycol. This does not adversely affect the chiller selection by more than 5 tons. PHB however will need to add some glycol to the system when it is refilled. Also pressure taps on the top of the building should be routed to a drum for collection. These are extra cost to the project that my have to be incorporated. TZ and TTC decided that keeping the electric heat tracing on the piping is still a good idea since it is already included in the project.

02/27/06 – Pre Construction minutes sent to contractors. PHB submittals arrived today. TEI did not select this chiller on glycol. TEI has to contact the manufacturer about derating the chiller for glycol and see what the final tonnage is for this building.

02/09/06 – Pre Construction meeting held.

12/13/05 – Contracts received, corrected and shipped back to the County.

12/01/05 – Trane called to tell everyone that prices were increasing this day and they want a confirmed order. TTC and Tony told Bernie at PHB that they could hold the unit for us while we wait for final contract approval. It was noted that the bid is still valid for 60 days so this price increase should not affect our project. Trane should be made aware.

11/04/05 – TTC, Tony and Bernie finalized bid numbers and some issues with was actually submitted to the County on bid day. Tony gave TE the okay to write the contract for this project. TE will have contract finished by Wed of next week.

10/24/05 – Final Addendum went out last Friday for project. Bid opening is set for this Thursday.

10/14/05 – TTC left a requirement about testing in the spec that has to be removed. Single compressors on each circuit are the only one that York makes so TTC will have to pick that up in the Addendum. The prebid meeting was successful.

10/04/05 – TE has finished the drawings and specs for the rebid and is sending it out for bidding to HamCo Purchasing.

10/03/05 – IBI confirmed today that with the 800A breaker on the 10th floor we can buy the 150-ton chiller as long as the MCA is less than 800A. Our 150 ton chillers are 733A so we can go with the full sized chiller. TTC and MS will finish up the drawings today and will deliver them to the County on Tuesday.

09/27/05 – Tony gave TE the okay to go with a 125-ton chiller if needed. In talking with George Mullins and Mike Stiles today, there may still be a change to stay with the 150-ton chiller because the total cost of the electrical project has decreased from a $105,000 job to a $25,000 project. With this additional money we can upgrade the GE chiller disconnect to 1200A. This will allow us to increase the chiller size back to 150-ton. Mike and George have to setup a field trip with Joe Merkt to verify some electrical data to make this happen. We are still trying to get the rebid done by this Friday.

09/18/05 – ThermalTech began re-engineering the project for buying the chillers on LCC and re-doing the electric design to meet the new standard for providing dual power to all equipment in the building. We want to be complete with the rebid this Friday but are running into some roadblocks where it concerns feeding the new rooftop chiller with an 800A max service or MOCP.

by TEI TEI

ADMIN – 7th Floor Engineer’s Office Renovation

08/31/07 – TEI was instructed to put project on hold till next year.

08/08/07 – Project estimated cost seems to be over budget. CRA awaiting direction from Mark Donnelly on scope/cost coordination.

07/30/07 – TE has moved to the design phase of this project.

06/29/07 – TE is awaiting drawings from Cole & Russell. When the drawings arrive TE will put together a cost estimate for the project and present it to Hamilton County Facilities Department.

06/22/07 – TE received word from Cole & Russell that this project is moving forward. TE is now awaiting drawings from Cole & Russell.

06/15/07 – TE is awaiting Joe Merkt, the building manager, to return from vacation so that we can perform the necessary field work to move this project forward.

by TEI TEI

ADMIN – Pat Thompson’s Office Renovation

07/26/07 – CRA instructed to put project on hold till next year.

06/29/07 – Pat Thompson is reviewing the drawings. TE awaiting word to proceed.

06/15/07 – Meeting was held with Mark Donnelly on June 8, 2007 for owner review. TE is proceeding with issuing this project for construction to Cole + Russell Architects.

by TEI TEI

ADMIN – Christian Sigmund’s Office Renovation

07/26/07 – CRA instructed to put project on hold till next year.

06/29/07 – Christian Sigmund is reviewing the drawings. TE awaiting word to proceed.

06/15/07 – Meeting was held with Mark Donnelly on June 8, 2007 for owner review. TE is proceeding with issuing this project for construction to Cole + Russell Architects.

by TEI TEI

ADMIN – ADA Title I & II Upgrades

12/20/07 – Approved final electrical permit has been received.

11/06/07 – Punchlist completed.

09/28/07 – Pre-Const meeting held 09/18. Signed contract & Notice to Proceed issued. CRA received and approved shop drawings.

09/19/07 – TTC emailed the IBI permit to MD with the paid balance recorded and the permit number.

09/05/07 – Nick Richmond wrote in an email to Mark Donnelly “We have completed the plan review for this project. Mike Huntwork will be dropping of payment for permit tomorrow. On his way back from IBI he will be stopping by your office and dropping off (3) full sets of the drawings for this project. (1)set is for you, (1) set is for Bert, and (1) set is for trades.”

08/06/07 – CRA received Prosecutor comments including comments on Supple Conditions.

07/27/07 – GC Contract (including Alternate #1) sent to Mark Donnelly.

07/18/07 – CRA submitted the Building Permit Application to the City of Cincinnati Buildings and Inspections.

07/09/07 – Bids received. Evaluation complete.

06/29/07 – This project has moved out to bid. Addendum # 01 has been issued.

06/08/07 – TE received drawings form Cole & Russell. This project has moved into the Design Phase. Drawings have been set up and TE has investigated where to get power for the new door openers. TE will perform a site visit early next week.

05/01/07 – 3 drawing packages authorized: signage, trade & bid. CRA issued for Signage and Trades. Existing door closers cannot be adjusted to required closer pressure so eliminated from project.

04/27/07 – TE awaiting drawings from Cole & Russell.

04/20/07 – TE still awaiting notice to proceed.

04/13/07 – TE still awaiting notice to proceed.

04/06/07 – TE still awaiting notice to proceed.

03/30/07 – Meeting was held to assess final requirements. Awaiting notice to proceed.

03/23/07 – Next meeting is scheduled for this coming Monday. Cost estimates will be provided by TE CRA.

03/20/07 – Kickoff meeting complete. CRA building study is being revisited for newer additional ADA. See CRA website.

03/14/07 – NR created project journal and filled in back data from previous meetings. Meeting minutes from the kickoff meeting held the week of March 5 have been issued to project team from Tom Kamphake today via email.

by TEI TEI

ADMIN – AUDITOR’S 4TH & 5TH FLOOR RENOVATION

04/07/08 – Pre-Award meetings are complete. CRA and TEI sent out copies of the meeting minutes too project members. Letters of recommendation are forthcoming from the partners. CRA is submitting for the HVAC and Building permits today. TEI is submitting for the IBI permit tomorrow. Kevin Maue is supposed to send TEI a release note on the Data Aire unit with Debra Kuempel. A pre const mtg should be scheduled in late April. KM stated that a Notice to Proceed letter should come out on or about May 8th.

03/28/08 – Bid Results are published. PDF

2/07/07 – CRA furniture and move services package ready for issue.

11/30/07 – TEI has the pre-purchase specs on the street for the UPS and the Computer Room AC Unit. These pre-purchase bids are scheduled to be opened on 12/5 at the HamCo Purchasing offices. We have decided to push back a design review meeting until after the equipment is purchased so final design will be exact.

10/30/07 – Meeting was held to review the proposed layouts. See Meeting Minutes here. The next meeting is scheduled for December 4, 2007 at 10:00 AM at 3rd Floor of C.A.B. Auditor’s Office.

10/09/07 – Meeting was held to review the proposed layouts and to discuss project construction phasing. HCFD stated that the project would bid around January 2008. This will allow construction to start around March 2008. See Meeting Minutes here.

10/01/07 – Meeting was held as scheduled. HCFD requested that C+RA, IPM and TEI issue a CD Owner Review Set on 10-05-07. See Meeting Minutes here.

09/28/07 – CRA has a coordination meeting with Mark Donnelly scheduled for 10/01/07.

09/19/07 – TEI had a coordination meeting with CRA on HVAC, Elec and Fire Protection. TEI has to update to CRA latest floor plans and then incorporate the furniture into the project. TEI is behind on finalizing the computer room layout with Stewart but TTC has scheduled a field trip to the space on Friday to go over HVAC changes, fire protection changes and electrical changes to the space. We also have to route chilled water line from the mechanical room to the auditor computer room. MD asked for drip pans under the city water unit and the liebert unit. TK sent a new sent of final floor plans over for our use. TTC gave everything to Chris and Mike to update.

07/26/07 – Restart meeting was held. CRA and IPM are working on re-program and concept layout. Target dates are: August 8 review meeting, September 3 out to Bid.

06/27/07 – Project is on hold but may recommence soon. Awaiting authorization to start.

06/20/06 – Project is on hold due to funding problems.

06/06/06 – CRA Schematic Design documents and estimate have been completed and approved.

05/22/06 – CRA received the signed SOSS.

04/17/06 – CRA submitted the SOSS.

04/07/06 – Kick-off meeting was held. Field verification and photographs were taken.

by TEI TEI

ADMIN – LOBBY DESK RELOCATION

12/07/07 – CRA sent contracts to BW on 11/29/07.

11/30/07 – The bids have been opened and although, TEI wasn’t at the bid opening, we believe HCG may have won the bid. CRA would have to confirm.

11/15/07 – CRA BW directed to prepare contracts for HGC.

11/12/07 – TTC and MH has sent out Addendum 01 through CRA. The bid opening is 11/15. Zach sent Addendum out for bidders. Addendum included information on relocating and interior light switch and alarm security panel by the EC while the owner has agreed to have the HVAC sensors moved by JCI or by their HVAC team.

10/30/07 – Pre-bid meeting held at C.A.B. first floor lobby at 1pm.

10/19/07 – Bid documents completed. Bid issuance date is 10/23/07.

09/25/07 – Cost estimate entered into Archibus.

09/14/07 – CRA received drawings from TEI. Drawings updated. Awaiting approval to complete and issue for permit.

08/28/07 – CRA issued DD drawings.

08/02/07 – Project being resurrected. Kickoff meeting was held.

by TEI TEI

ADMIN – Welcome Desk Relocation

12/31/09 – This project is on hold per Bert Watts. It is archived on 12/31.

07/29/09 – SOSS have been submitted but not approved. CRA turned over schematic design drawings and cost estimates for review by the County. Bert has to present these to Jeff Aluotto for approval before final design can begin.

by TEI TEI

ADMIN -Auditor’s Office – 4th & 5th Floor Renovation

02/08/10 – Payapp has arrived. TEI finished the as-builts drawings. We are shipping the original record markup dwgs, cd’s and final payapp to Bert for closeout.

02/04/10 – Look i can’t take it any longer. This project has to be moved to the Archive. It is just too old. I will update the final QCE payapp when it arrives. It is being moved today.

01/28/09 – TEI contacted QCE but Cavanaugh is out. We have not received their final payapp. As-Builts are almost done with Cad markups by TEI.

12/31/09 -QCE turned in closeout documents today for approval to TEI. TTC received and approved. As-Builts will be updated in 2 weeks by TEI. We need a final payapp from Queen City still for approval which is coming next week according to Mike C. After that this job will be closed out.

09/22/09 – We have contacted Queen City several times about closeout and still have not received the documents.

08/20/09 – QCE payapp has been approved but we have not received their closeout documents yet.

07/29/09 – QCE reported today that they know the O&M’s are due but they have just been too busy on trying to get new work in today’s market to finish them up. TTC told them that they could bill for retainage on this project and leave the closeout document line item at 0% until they submit the O&M’s. Cavanaugh said he might do this but they will work on the closeout docs.

06/23/09 – The final EDI item is the revised sequence which TEI wrote and sent to Data Aire and EDI for approval. The sequence has been approved but TTC is waiting in equipment deadband from DA. This is not critical to finishing the sequence. TTC may just update and send final to Bert. All other project contractors are complete. Get an update on closeout from Bert on EDI, QCE, CDFP paperwork.

05/22/09 – Data Aire Unit is running as designed on both chilled water and DX cooling. The system is changing over quickly and keep the space cool. TEI to write final sequence for the project.

04/15/09 – CR Architecture and Design awaiting final punchlist items to be completed.

04/03/09 – All work and punchlist items have been completed by EDI on this project. O&M’s have been logged in. Final payapp has been received. The payapp #2 was on hold but has now been release to Bert. The CDFP closeout paperwork was signed and sent to Bert yesterday. Queen City still owes us final payapp and O&M’s on this project. EDI still has to use chillers in the building before we know everything works.

02/18/09 – A controls meeting is scheduled for Friday with EDI and Data Aire.

02/07/09 – TTC requested a revised payment from EDI for this project.

02/02/09 – Terry & Mike will be down for project punchlist on Feb 4 for final walk through. The new insulation worked perfect through the cold snap last week. Bert reported no condensation on the unit, even on the the through metal parts.

01/22/09 – EDI is at the site today installing the insulation inside the unit today. All interior surfaces.

12/20/08 – Problems with the Data Aire Unit still exist. Condensation is a real problem. The unit is not insulated. DA said they can fix this problem with insulation. The controls and the chilled water pump are still not interlocked. A JCI N2 card can help the situation. Duft said that would be the easiest to use. TTC said it was going to be provided at no cost from DA on this project.

11/25/08 – TEI is working feverishly with EDI and QCE for final change orders. Hours were spent trying to resolve QCE change order #3 today and it still wasn’t exactly right. Another $317 needs to be written for QCE. ThermalTech must complete final punchlist on this project.

11/24/08 – Project is 95% complete. Final Change Orders need to be in by 11/26. TEI is working on them today.

10/16/08 – Bert reports that Stewart will be making all final computer moves this weekend and by Monday all essential Auditor equipment will be on the 5th floor. Demo of the 4th could start shortly for all contractors. QCE has made more modifications to the plug strip on the server racks and disconnected the 4th floor rack to facilitate its move to the 5th floor. The project is finally making progress towards completion.

09/18/08 – Trying to setup TCO for next week.

09/17/08 – Server must be running by next Tuesday. Fan test not complete. Temporary route condensate to 4th floor hub drain. EDI ductwork not installed. Power outage has delayed final fabrication of ductwork.

09/04/08 – Elevator quote came into today for $3k. Work to be done next Monday with Thyssen Krupp and EDI.

08/2508 – TTC sent out the plumbing Bulletin drawings to Helmes plumbing et al.

08/22/08 – TTC visited the site today for floor drain relocation. UPS is in and being wired. Disconnects are installed. Fire protection is in 5th floor room.

08/08/08 – Const Mtg notes from today include

08/01/08 – GM located new locations for the bus plugs disconnects for the project. One on the 5th and one on the 4th. QCE will have to move an existing disconnect to make the new ones fit.

07/30/08 – Data Aire Unit ship date is August 20. It will arrive 3 or 4 days after that. Quick ship prices were given but overly costly to the project for the minimal gain.

07/17/08 – Const Mtg. We discussed project schedule for this project. Contractors reviewed progress. Demo found a stairwell pressurization fan in the new computer room that must remain as is. EDI will seal ductwork. Computer room unit delivery is 4 to 5 weeks.

05/30/08 – Pre Const Mtg complete. Equip 5 weeks out. Contractors to provide schedule for start of the project.

by TEI TEI

ADMINISTRATION 8TH Flr Planning Renovation

08/13/10 – The space looks great. Tenants say they are not used to the new work flow yet but will give it time.

07/23/10- -Work continues. BW is handling the Const Admin of the trades work on this project. No new questions or issues that TEI is aware of.

05/26/10 – TTC will check on progress while at the permit office on Thursday.

05/13/10 – City sent permit receipt to TEI for use on this project. All permits look to be approved at the City Web portal.

05/05/10 – Check on progress with Dave S and Bert W. IBI inspections for rough in are complete. Cindy office is under nightly construction.

04/26/10 – Work has begun. Phasing was taken over by Bert. We check the scale of the drawings on D size prints and it was right, check with Bert on Thursday to see what the issue is.

03/24/10 – Arch permit review is complete. The changes will take some time and we need updated specs from Speer. These should be permitted on Thursday/Friday.

03/23/10 – Permit drawings are being reviewed inside ThermalTech and at the Architects office. We should be acquiring permits by Wed or Thurs this week. We have incorporated all the client comments into the drawings and added specs to the job.

02/24/10 – We had the SD review meeting. Bert asked for a new SOSS for permitting the job and incorporating the changes and Archibus input. Smorey requested a new schedule. TTC will get that completed on Friday.

02/04/10 – We are still trying to get an SOSS worked out with Public Works for this job and Bert is assisting us in this attempt.

01/28/10 – TTC and Bert discussed a new SOSS for County Trades installation and TTC provided it. There is some doubt about how will do this work and how it will get design. TEI will wait for Facility guidance before proceeding.

12/21/09 – TTC, BW and Smorey met on site to go over the changes required in the renovation. TTC got drawings copied for Hamco and TEI and is preparing the SOSS. This project will require electrical, fire alarm, mechanical and general work to complete.

by TEI TEI

ADMINISTRATION – 7TH Flr Engr Furniture & Electrical

03/30/11 – Project is complete. Art Smith & TJ Williams invoices has been sent to Pat Day for final payment.

02/22/10 – The initial stage is complete. I believe I am waiting on the County to find some furniture pieces for us to use to fix up the remaining parts. I haven’t heard back from Judy on how well the repairs went.

01/24/11 – We have finished the first pass through on the furniture fix up. I need furniture pieces to have Art Smith fix up the remaining parts to get it as good as it gets for its age. I would like to move to next phase and replace the front counter after we get our contract and counter SOSS approved.

10/25/10 – TJ Williams bid was resent to Judith. Engr needed more information. Ginter has to update their bid and we will forward that along.

09/29/10 – Bids were collected and sent to Bert. The 2 day repair cost is under $2000.

07/28/10 – Need Judith contact info. George Mullins may have to design this project with Susan.

07/23/10 – ThermalTech attempted setup a design meeting for SD and Cost. Judy email bounced back so I need to find her new number.

05/26/10 – SOSS approved. We will setup fieldwork and kickoff next week.

04/26/10 – Schematic Drawing is complete. Fieldwork is complete for SD. We are working on cost estimate to send with SD.

03/23/10 – Project on hold.

01/26/10 – Bert contacted ThermalTech about permit requirements and was told the mechanical code would require a permit for the new ventilation calc required in a new conference room. Bert has the SOSS for this project.

12/31/09 – Bert contact TEI to work on this project and renovation. TEi owes Bert an SOSS based on the HamCo Scope of Work.

by TEI TEI

CAB – Access Control Project

09/22/10 – SOSS Approved. Review comments sent and received from Bert.

09/29/10 – Markups corrected from BW. QC check complete. TEI needs to pick up NV marks and send to Purchasing on 10/1.

10/28/10 – Contracts for ICS have been sent to Bert Watts for approval. Project is under budget. No allowance has been added because of a bonding issue.

01/24/11 – Bert and George decided a pre const mtg will be set next week to kick off this meeting.

02/22/10 – Paul Wilson has made good progress on this project. Wiring is ran on upper floors and they are working their way down to the dock.

03/30/11 – Paul Wilson submitted and change order to remedy the power supply issues on the tenth floor. REX needs to be changed so the door does not open at night when approached by security. GM said that the contractor can reprogram this feature. Project seems on track.

04/27/11 – Bert reports that installation is very nearly complete and the system is ready to begin use this week or next.

05/13/11 – Punch list was issued to IPS. Project should be wrapping up.

06/29/11 – IPS has submitted its main pay app, now we need to close out the rest of the project.

08/28/11 – New Signed change order came back. IPS should be doing this work soon, rest of project complete. We can request closeout at any time.

09/28/11 – All work to be completed has been put on hold per TTC.

10/24/11 – The certified payroll for IPS for the dates 10-2 through 10-9 was sent to Hamilton County and TEI.

11/15/11 – IPS requested final payment from Hamilton County in the amount of $4,984.22. TTC responded by saying all paperwork had been sent to Hamilton Co. and IPS should get paid per the Contract.

12/23/11 – This project will be archived for 2012.

by TEI TEI

CAB – Sheriff’s Re-entry

01/20/12 – Mike B. spoke with Mr. Gaylord Poe, President and Chief Electrical Inspector for IBI regarding the CAB space. Mike explained that there is minimal electrical work being completed on the project, and that the electrical connections will be able to re-use the existing circuits that were removed from the systems furniture when it was disassembled. In order to get a permit, a plan exam is required with IBI. Gaylord and I discussed that the net electrical load on the space will be significantly lower than what was there previously. Gaylord said that if we draft up a letter, and submit it to him directly, he will approve that as a plans examination, and we will be able to then secure the permit. Mike has drafter the letter, and will run the letter with the permit fee to IBI asap. Actual plans, panelboard schedules, single line diagrams, etc will not be produced for this project.

01/30/12 – Mike walked through the electrical permit at IBI, and then dropped off paperwork to Steve Wessels. A final electrical inspection has been requested for Wednesday Feb. 1st. Upon verification of the inspection, Mike will send a follow up e-mail.

03/07/12- Walk-through permit was conducted by GBBN. Permit approved conditionally requiring Mechanical Permit to validate movement of existing linear diffusers to other side of new partition.

05/30/12- Bert Watts requests As-Built drawings to be completed. He indicates the Building Department has given final approval of construction.

by TEI TEI

CAB – First Floor Security Cameras

02/22/12 – SOSS Approved. TEI and Hamco to have a kickoff meeting for the project on 02/23/12.
03/27/12 – On 03/09/12, ThermalTech issued drawings for County Review. Received comments from Bert Watts and Chris Neff, waiting on final comments from Rodney Lofland to complete project, and issue and ITB.

5/30/12 – ThermalTech submitted drawings and specifications to County for final review (with all county comments incorporated). Waiting on notification from county to proceed with issuing for bid.

6/26/12 – Project on hold until notified otherwise from BW.

by TEI TEI

CAB- 7th Floor Engineering Office- Phase 2 Renovations

06/15/11 – SOSS delivered to Bert Watts.

07/08/11 – GBBN informed the County does not have the required funds to proceed with project at this time.

07/20/11- Meeting between design team and Owner/User Group.

03/20/12 – SOSS re-submitted to Bert Watts.

03/26/12- Bert Watts notifies GBBN that Judy Sellmeyer has retired. Her replacement is trying to get up to speed on projects.

by TEI TEI

CAB Beam Repair

Partner: THP

03/20/13 – THP met with County to review beam condition in the basement of the Administration building. Based on the initial visit an additional assessment needs to occur to determine the effect of the active deterioration on the beam. The assessment will determine the need to stabilize or repair the beam.

04/04/13 – THP and County reviewed available drawings.

04/17/13 – THP issued SOSS.

04/22/13 – County declined SOSS due to funding. THP published closeout email; answering County questions.

by TEI TEI

ADMIN Boiler Flue Phase 2

Partner Lead: TEI

09/26/13 – At Partner Meeting we discussed including as an Alternate to my rebid package, adding some additional tuck pointing. TTC to coordinate this with Tony Matre.

10/01/13 – THP issued SOSS.

10/09/13 – Hamilton issued executed SOSS for THP.

10/21/13 – Bid Package out for Bid.

10/31/13 – Pre-Bid schedule for 11/01/13.

01/21/14 – Tony said project is not scheduled to start yet. It is probably on hold.

by TEI TEI

ADMIN Public Defenders Renovation 9/10

Partner Lead: GBBN

08/28/13 – GBBN and ThermalTech met with the Public Defenders about possible office relocation into the Admin bldg on the 9th or 10th floor. PDO indicated they preferred having two departments on separate floors; one for guardians and one for juvenile. On either option ThermalTech believes they can shrink the computer room down to a smaller footprint and make it work in either option. ThermalTech went back to the computer room and did a take-off of all existing equipment for a more complete layout the new computer room to schedule a meeting with HamCo CMS soon.

08/28/13 – TEI sends markups of Roof Top Equipment affected by scope of work.

08/30/13 – Facilities indicates the 10th Floor scope will not be approved.

08/31/13 – GBBN/IPM send floor plan layout indicating the 10th floor CLEAR and Community Development areas together would be suitable for receiving the GAL department to Facilities.

09/02/13 – Facilities contacts GBBN to inform that the 10th Floor scope will be reconsidered.

09/06/13 – Facilities requests a revised SOSS be submitted for the Tenth Floor relocation scope.

09/06/13 – GBBN submits revised SOSS to Bert Watts for processing.

09/10/13 – Diana Supe requests dates of availability for Kick-off meeting with Public Defenders Office and Facilities Department.

09/11/13 – GBBN received approved SOSS from Bert Watts.

09/17/13 – Re-kick-off Meeting held at 230 WHT Law Center with PDO, Facilities, GBBN, IPM and ThermalTech.

09/25/13 – Team met with PDO and Facilities to discuss computer room options in center of building and present initial cost estimates on all 3 options for 9th & 10th floor work. Everyone agreed to the computer room being built around the 4 communication racks, modifying the waiting area directly off the elevators, putting attorney offices on the west wall and leaving a growth area to the west of the computer room.

09/26/13 – At the partner meeting we learned that there may still be VAT tile remaining on the 10th floor CLEAR area and the 9th floor offices outside the computer room. GBBN will update the cost estimate for these options.

by TEI TEI

ADMIN 610 DOC Room

Partner Lead: ThermalTech

03/04/13 – SOSS submitted to Bert.

04/24/13 – SOSS approved and PO issued to ThermalTech. ThermalTech to schedule Kickoff meeting with Bert Watts to establish schedule.

6/26/13 – TEI has sent drawings to HamCo for review. Need to discuss HamCo comments and issue for bids.

08/28/13 – Pre-Bid meeting conducted on 8/28/13. It was observed that (2) fiber optic data cables were already installed from the data rack in the closet by the freight elevator to the A/V cabinet in Room 610.

09/25/13 – Bids received, Phasor Electric is Low Bid. Contracts are into county for approval.

10/30/13 – Phasor to begin work week of November 4.

01/24/14 – Work completed. Phasor to submit final as-built drawings and project can be closed out.

02/25/14 – TEI reviewing as-built drawings. All other close out documents have been turned over to Hamilton County.

by TEI TEI

ADMIN Boiler Stack Removal

Partner Lead: ThermalTech

05/07/13 – Met Geiler on site to discuss steel stack issues.

05/21/13 – Met Van Packer & Geiler on site to discuss removal of the Van Packer stack vs the standard 1/4″ steel stack. We asked Geiler for a change order to remove the entire stack. Regardless of what we do with the existing stack we told Geiler that we wanted to do the EGEN exhaust stack work in this phase.

06/25/13 – Geiler cannot find a contractor willing to take down the stack. ThermalTech asked for a 2 week schedule to put the EGEN exhaust modification in for the base bid work. We will have to address the remaining funds as a deduct change order.

08/19/13 – We had Pinnacle check the EGEN stack and muffler for asbestos. All samples came back clean. Geiler required that report in writing. Tony Matre provided to contractor.

08/26/13 – Coredrills are complete. EGEN exhaust has been rerouted through outside wall. SS end cap and birdscreen is installed. 10″ sleeve is installed but needs to be trimmed back. Insulation needs to be reinstalled. ThermalTech instructed Geiler to get a change order price from R&P in Lexington to do the work. We will decide on a change order after Tony sees how much money is left in budget at end of fiscal year 2013.

09/26/13 – ThermalTech informed Geiler to wrap up EGEN work and provide a credit for the boiler stack work. This scope will be rebid this year.

10/01/13 – THP issued SOSS.

10/09/13 – Hamilton issued executed SOSS.

10/18/13 – THP issued drawings to TTE.

10/29/13 – TEI sent out back punchlist items to Geiler and asked for project closeout.

01/24/14 – TTC contacted Geiler for final closeout and gave them a deadline of 1/27 at 4pm for closeout documents.

02/24/14 – Project still not closed out. TTC will contact Geiler again.

03/21/14 – Geiler came down to finish the insulation and was held up by someone. Tony said it was the County that stopped them. TTC asked Doug to find out who stopped them but I have not heard back yet.

04/04/14 – Geiler has come back two more times to work on insulation and replace one sprinkler head. We are still waiting for final closeout and change order documents.

05/22/14 – Punchlist completed. Waiting on County approval to closeout.

06/02/14 – Geiler contacted again for closeout docs. Doug Weberding says it will be soon.

07/30/14 – Geiler sent closeout docs to Tony but Payapp and change order are both wrong. TTC sent revised Payapp and new change order to Geiler for signatures.

08/15/14 – Geiler inquired about final payment. TTC forwarded to Tony.

by TEI TEI

ADMIN Energy Star 2014

Project Lead: TTE

09/23/14 – SOSS approved. TTE to setup verification field trip with Rich.

11/28/14 – The EPA sent several questions in about the Data Center and TTC answered them immediately with pictures, graphs and charts explaining the validity of this data center.

12/12/14 – Building still listed as “Under Review”.

12/18/14 – Energy Star was awarded to this building.

by TEI TEI

ADMIN Facade Repairs

Partner Lead: THP

08/29/13 – County requested SOSS to provide repairs to the Terracotta at the first and second floors.

09/09/13 – THP met with Mike Florimonte to discuss scope and review the conditions.

09/13/13 – THP issued SOSS and Tony Matre verbally authorized work to begin.

09/23/13 – THP completed field survey.

09/25/13 – THP completed rappelling survey.

10/02/13 – THP issued Owner Review documents.

10/10/13 – Bid Documents Issued.

10/18/13 – Prebid Conference.

10/24/13 – Addendum #1 Issued.

10/30/13 – Bid Opening completed with five contractors under the budget. County reviewing bid proposals.

11/05/13 – ZSR Construction was the low bidder.

11/11/13 – THP sent AIA contracts to the County for review.

02/20/14 – County returned marked up Contract to THP for final publication.

02/21/14 – THP sent updated Contract to ZSR.

02/25/14 – ZSR sent completed contract and forms to THP.

02/25/14 – THP sent final contract documents to County for execution.

03/27/14 – ZSR has received an executed contract. ZSR and County need to schedule the start up.

04/01/14 – ZSR submitted submittals to THP.

04/07/14 – THP returned submittals to ZSR.

04/10/14 – ZSR submitted a construction schedule. Mobilization and work is planned to start 6/2/14.

05/29/14 – Waiting on project to start.

06/16/14 – ZSR started work. Terra cotta and concrete repairs in progress on West and South elevations. Work is occurring at night to keep entrances open and minimize noise and dust interruptions. Expected completion date is end of July.

07/31/14 – ZSR completed all repair work. ZSR working on coating work.

08/15/14 – Construction is complete.

08/18/14 – Pay Applications 1 and 2 delivered to Tony Matre.

08/19/14 – Concrete patch included paint to match surrounding area. THP rejected the patch application as not matching. ZSR completed a second paint application and Tim Bennett noted the match was acceptable to THP. THP recommend, County review the color match prior to accepting the coating color.

09/25/14 – ZSR completed all repair work. THP awaiting warranty/closeout paper work and final pay application.

10/13/14 – ZSR submitted all closeout warranty paper work and final pay applications to THP. THP forwarded approved closeout paper work to the County.

10/16/14 – Closeout paper work and final pay application submitted.

01/29/15 – Project complete.

by TEI TEI

ADMIN Recorder’s Storage Room

Partner Lead: GBBN

01/24/14 – Project Scope meeting scheduled for 01/28/14.

02/03/14 – THP – GBBN requested possible structural analysis of storage relocation project and THP provided an hourly estimate for THP services.

02/25/14 – GBBN and TEI met with Bert Watts to investigate the possibility of relocating the file storage to the CAB 10th Floor CLEAR area. GBBN generates a preliminary sketch indicating the files will fit into the area designated at the meeting. Bert Watts indicated General Services will be used to generate a Schematic Design and Opinion of Probable Cost for this scheme.

02/26/14 – GBBN confirms a second access door will be provided into the File Storage and the adjacent workstation will be removed to allow for the door placement. Revised cost estimate is due this Friday to Bert. ThermalTech believes this work can be done without an additional Liebert cooling unit given the HVAC design and CFM’s currently in place in this space.

03/12/14 – GBBN issues SOSS to Bert Watts involving change to 9th Floor Relocation in the existing CLEAR offices.

04/23/14 – GBBN, ThermalTech and Bert kicked off the design of this project today. Design drawings are due June 26 partner meeting. Work to begin after CLEAR moves before December.

03/13/14 – THP noted to GBBN an initial analysis of installing shelving on the ninth floor is possible and a full study will be needed to confirm the proposed final loading. THP based the recommendation on an earlier study in 2012 for a proposed shelving arrangement.

04/24/14 – THP – GBBN approved the full analysis.

05/19/14 – GBBN receives dead load figures from Republic Storage for the existing newer shelving units. This information is forwarded to THP for their load analysis.

05/27/14 – THP completed the analysis of the proposed shelving solution on the ninth floor. THP does not recommend the proposed loading configuration, based on the current information and proposed arrangement. The 2014 shelving arrangement substantially increased the loading from 2012 arrangement. To allow shelving to be moved to the ninth floor, THP recommends either spreading load over additional bays or investigate the slab to determine the slabs reinforcing.

05/27/14 – THP forwards Load Analysis report to GBBN indicating the existing floor loading capacity will not support the current layout for relocating the existing file storage. ThermalTech started design and created floor plans, put on hold for load analysis.

06/10/14 – THP – Terracon completed scans of floor to determine bar sizes.

06/14/14 – THP – Terracon published report to THP.

06/24/14 – THP – THP completed an analysis based on the Terracon report and published a letter to the County for review.

07/31/14 – THP study complete. Bert says project funds are being reallocated, this project pseudo on hold.

08/28/14 – THP – Correction to 7/31/14 entry. THP portion of study complete.

by TEI TEI

ADMIN Heroin Task Force

Partner Lead: GBBN

04/08/15 – Kick-off Meeting with Bert Watts to review site and discuss scope.

04/08/15 – AFSS scope clarified with ThermalTech related to placement of new wall related to existing sprinkler heads.

04/10/15 – GBBN sends door hardware schedule to Facilities for purchase of door and frame. GBBN coordinates electrical scope for door hardware access control with ThermalTech.

04/15/15 – GBBN and ThermalTech provided notes and code interpretations for the owner to self-perform work on the existing spaces to fast occupancy. We will continue to support the owner as permit and construction questions arise.

04/17/15 – Drawings were issued for permit which include General Architectural and Electrical Drawings. GBBN attempts a walk-through permit. Minor door hardware corrections requested related to access control code section 1008.1.4.4.

04/21/15 – GBBN revises drawings and resubmits to City of Cincinnati Buildings and Inspections. Plans Examiner requires change to owner requested pull station for access control hardware.

04/22/15 – GBBN revises drawings and resubmits to City of Cincinnati Buildings and Inspections and obtains Building Permit during walk-through inspection. Building Permit delivered to Facilities Department. Building Permit #2015P02980.

05/26/15 – TEI picked up and dropped off to Facilities Department the Electrical Permit from IBI, permit Number CIC1501884.

07/22/15 – Install substantially complete, and TEI drafted a letter to John Nester with electrical loading information outlining the field changes that occurred and the effects that the field changes had on the service calculations. John Nester stated that trades electricians will submit to IBI on final inspection.

08/27/15 – JCI extending relay to door in space for fire alarm interface.

09/24/15 – At Partner Meeting, we said project was pretty much complete. Is it ready to archive?

10/28/15 – Project read to archive?

by TEI TEI

ADMIN Re-Entry to CLEAR Space

Partner Lead: TEI

01/07/15 – Bert Watts contacts GBBN about Kick-off Meeting.

01/13/15 – GBBN and TEI visit site with Bert Watts for kick-off meeting. After meeting both firms send SOSS to Facilities Department.

01/14/15 – GBBN sends updated schematic design drawings to Bert Watts and TEI for review. TEI initiates Mechanical scope drawings to each.

01/23/15 – GBBN requested by Facilities Department to revise drawings for the Auditor proposed relocation to CAB First Floor.

01/27/15 – GBBN, issues revised preliminary floor plan of Auditor’s proposed relocation to CAB First Floor. Facilities sends to Auditor for review.

03/18/15 – ThermalTech sent in revised SOSS for Architectural and Engineering estimates for approval. We kept the AE fees under $50,000 but that makes the contingency left on this project lower. The amount of system furniture that can be used and is in storage is in question.

04/07/15 – Project is sort of on hold. Some work is being done as a basic necessity now but as the final details are flushed out the partners are on hold.

04/30/15 – BW requested new SOSS for moving Sheriff Reentry to 10th floor in much the same configuration as 9th. GBBN to have TTE under their SOSS, meet with Reentry program and try to save Facilities storage space on the same floor.

05/19/15 – GBBN requests clarification of scope for original SOSS.

05/27/15 – Bert Watts requests SOSS be revised in scope to only included Schematic Design and Estimate.

06/25/15 – Bert said that CLEAR moved out this week and this project will probably start in July or August.

08/10/15 – GBBN requests clarifications from Facilities on scope of work. Facilities responds with associated answers. Facilities issues Archibus Pop-up Project Request Number for estimating input.

08/19/15 – GBBN completes Architectural scope estimate in Archibus.

08/20/15 – TEI completes Mechanical and Fire Protection scope estimate in Archibus.

08/25/15 – TEI completes Electrical scope estimate in Archibus.

10/27/15 – GBBN sends revised floor plan to Bert Watts for review.

10/28/15 – Bert Watts indicates the revisions are acceptable. He will submit to Re-Entry for approval.

by TEI TEI

ADMIN Sheriff’s Fugitive Detective Relocation to 9/10

Partner Lead: GBBN

10/13/16 – GBBN, ThermalTech and Facilities Department meet to propose a new SOSS for two options for relocating the Sheriffs’ Fugitive Department to the CAB 9th or 10th floors. Facilities provides sketches for each option.

10/13/16 – ThermalTech sends preliminary labor costs to GBBN for the SOSS.

10/17/16 – GBBN contacts IPM about the project and obtains estimates for labor related to systems furniture and move services.

12/28/16 – GBBN finishes estimates and checks in with ThermalTech for progress. TTC for finish fist couple of days of 2017 after everyone returns to work.

01/23/17 – All cost estimates are updated in Archibus and were sent to Commissioners by Bert for funding determination for Option #1. Option #2 was removed because of the Halon cost associated with project.

by TEI TEI

ADMIN – Lobby Lighting/Finishes/Auto Title Door

Partner Lead: GBBN/IPM

07/23/15 – Bert Watts contacts GBBN about Kick-off Meeting. GBBN reviews attached documents and requests clarifications.

10/28/15 – GBBN and IPM meet with King Lighting to review scope of work at site.

11/03/15 – GBBN and IPM meet with King Lighting to place sample light fixtures on columns at site. Summary of discussions and decisions sent to Facilities.

11/04/15 – Cost Estimate for Auto Title Door project sent to Facilities.

11/05/15 – GBBN consults Facilities about lighting level issues at north elevator bank verses south elevator bank.

12/07/15 – King Lighting sends final cost estimate and fixture cuts to IPM for lighting upgrade recommendations.

12/07/15 – GBBN contacts directory vendor for estimates on directories for Elevator Lobbies.

12/08/15 – ThermalTech delivered cost estimate for installing extra power and installation for new light fixtures.

12/09/15 – GBBN and IPM to meet to discuss rendering views required and all cost estimate data received.

04/07/16 – Facilities requests GBBN author an AIA Change Order to KFI for the secured employee entrance inside the BMV vestibule.

04/27/16 – GBBN authors Change Order for new Secured Employee Entrance.

05/13/16 – GBBN receives authorized copy of Change Order.

by TEI TEI

ADMIN – 6th Floor Chief of Staff Renovation

Partner Lead: GBBN

11/30/16 – GBBN issues SOSS to Facilities.

12/15/16 – GBBN receives approved SOSS from Facilities.

12/6/16 – GBBN visits site with Facilities and photos and takes measurements for field verification.

12/14/16 – Drawings issued for permitting.

12/15/16 – GBBN completes drawings for owner review. GBBN receives drawings and Mechanical Permit application from TEI. GBBN performs new City of Cincinnati Building Department Same Day Permit Walk Through processes. Permit and approved permit is delivered to Facilities.

12/27/16 – Terry and John meet on site to discuss extra sprinkler location. TTC states that sprinkler is not need given the actual wall layout, it is exactly 7′-6″ from interior wall. This should save some money. If city requires and ECO, TTE and GBBN will take care of this at final inspection.

01/25/17 – IBI final electrical inspection passed, and tag placed on serving panelboard.

01/27/17 – Project is complete, waiting on the closeout documents from everyone. Ready to due as-builts.

02/01/17 – Revised energy impact study sent to Bert Watts for review. Small savings but ThermalTech used HAP to generate the savings and it matched the original estimate.

02/23/17 – John to send ThermalTech as-built markups for this job.

by TEI TEI

ADMIN – 5th Floor Roof Access Hatch

Partner Lead: THP

09/07/16 – Hamilton County requested THP review a proposal to add an access hatch and railings to the 5th Floor mechanical space.

09/26/16 – THP and County setup planning meeting to review conditions and proposal.

10/27/16 – THP reaching out to LGB to develop a proposal.

12/05/16 – LJB provided fall protection design services to THP.

12/06/16 – Draft SOSS provided for review. Hamilton County placed SOSS on hold.

03/10/17 – Tony Matre calls GBBN to discuss permitting issues and logistics concerning Safety Railing products used at Alms and Doepke Roofing Restoration project.

by TEI TEI

County Administration – BMV Cooling Issues

08/03/16 – TTC was on site with John Gilbert yesterday troubleshooting the BMV area. It looks like its fed from AH-1 not AH-3 as originally thought. County to investigate and get back to TEI. Airflow on CV box was also locked in at 20% and several boxes were set to 0%, I would assume due to overcooling.

09/25/17 – No movement on this project. TTC sent an email to Tony about this project and helping if needed.

by TEI TEI

CAB Heroin Task Force Security System

Partner Lead: TEI

08/16/16 – TEI meets with Lt. F of the Heroin Task force, and walks the project to develop scope of the project. TEI submits SOSS to HamCo facilities department. Project pending SOSS approval and PO.

09/28/16 – Project pending SOSS approval and PO.

10/25/16 – project pending SOSS approval and PO – per facilities, pending grant to be issued after first of the year.

01/25/17 – Project pending SOSS approval and PO.

02/22/17 – Project pending SOSS approval and PO.

03/22/17 – Project pending SOSS approval and PO.

04/26/17 – Project pending SOSS approval and PO

05/22/17 – Project pending SOSS approval and PO.

06/27/17 – Project pending SOSS approval and PO.

07/25/17 – Project pending SOSS approval and PO.

08/28/17 – Project pending SOSS approval and PO.

09/26/17 – Project pending SOSS approval and PO.

10/25/17 – Project pending SOSS approval and PO.

11/29/17 – Project pending SOSS approval and PO.

by GBBN GBBN

CAB – Conference Room #607A/B – Two Office Renovation

Partner Lead: GBBN
09/06/17 – Facilities contacts GBBN and requests that a new SOSS be authored for this first phase of the overall project. Facilities confirms the west wall of the existing Conference Room is to be converted to a glass storefront similar to the GBBN Standards Build-out proposal.
09/07/17 – GBBN issues SOSS to Facilities.

by GBBN GBBN

CAB – Sheriff Re-Entry Renovation Study

06/05/17 – Facilities contacts GBBN about new project and scheduling for Kick-off Meeting.
06/06/17 – GBBN sends Facilities and TEI quick floor plan of the scope discussed in the meeting.
06/08/17 – GBBN quick floor plan of the scope discussed in the meeting to IPM for further investigation. IPM agreed little other options are possible due to the tight spatial constraints. Any other options would not affect estimating.
06/26/17 – GBBN meets with Re-Entry Staff to discuss 2018 Project Request Estimates #1877/8 and possible differences from Estimate #1941.
06/27/17 – GBBN , Facilities meets with Re-Entry Staff to show Four Options for the Renovation. Option #3, without Interview Rooms was selected. Estimate #1877 for the 2018 Estimates was eliminated. Option #4 presented in the meeting will be used for the Estimating Purposes.
07/14/17 – Facilities reports back to GBBN indicating which final version of plan is to be finalized in the Archibus Project Request Estimate. TEI confirms the final layout will not dictate any MEP cost revisions to the estimate. The estimate has been finalized.
02/22/18 – Project ready to archive.

by GBBN GBBN

CAB – Conference Room #607A/B Office Renovations/CAB 6th Floor West Office Renovations/Standards

Partner Lead: GBBN
02/16/17 – Facilities contacts GBBN about new project and scheduling for Kick-off Meeting.
02/21/17 – GBBN meets Facilities and Traci Paquette at the CAB to hold Kick-off Meeting.
02/22/17 – GBBN contacts TEI to discuss scope of project and to discuss questions for Facilities concerning the CAFM/As-Built requirements. GBBN sends question to Facilities for answers. Facilities sends answers to GBBN.
04/26/17 – GBBN meets with Jeff Aluotto and Facilities to discuss expanded scope of work, for more of the same office suite. New finish palates, layouts and associated renderings will be generated. Various options for new office suite layouts and material selections for most of the west half of sixth floor.
06/09/17 – GBBN meets with Administration to present conceptual ideas for proceeding on the CAB 6th Floor West Office Suite.
07/20/17 – GBBN meets internally to discuss the latest developments on the project. GBBN later sends scope questions related to next steps and future regular status update meetings. Facilities responds to GBBN’s questions about next steps with the project and next SOSS. Preliminary contact with made with Theresa before Jeff Aluotto’s pending vacation. Facilities will report back with more input from Theresa concerning the scope required for the next SOSS.
07/26/17 – GBBN finalizes details with RCF concerning Mock-Ups and systems furniture. carpet and glass wall installation capabilities by RCF. GBBN forwards electronic drawing files to RCP for use in organizing mock-ups.
08/03/17 – Facilities confirms Trades can install Electrical Whips for new systems furniture.
08/08/31 – GBBN receives communications from River City Furniture Representative for State Term Contracting.
08/10/17 – GBBN authors SOSS, specifically for this modified scope of work, and delivers to Facilities.
08/25/17 – GBBN receives call from Facilities confirming project is on hold. Facilities requests construction estimate for scope of work be completed before GBBN terminates work.
08/31/17 – GBBN receives clarifications on scope of estimating from Facilities.

10/23/18 – Jeff Aluotto makes request through Steve Kenat that all latest Design information from this project be forwarded to his attention.

by TEI TEI

CAB – Chiller and Cooling Tower Replacement

11/17/16 – THP – Hamilton County requested SOSS from THP and TEI about this same time.
01/25/17 – ThermalTech receives signed SOSS and PO for this project.
02/02/17 – Getting data back from Trane, Carrier and York. No one can do Centrifugal at 208V for this job except Carrier but they have to remote mount the starters and controllers somewhere else. Trane.
02/16/17 – TEI, facilities and building manager meet to kick off project.
02/22/17 – ThermalTech working on LCC and marginal rates from Duke. Amy, Annie and Terry working on LCC worksheets. TTC got new Duke Rate sheet from Michelle
02/23/17 – The County provide direction today to size the new chillers for the addition of 9th and 10th floor in the future. check piping sizes and pumps and cooling tower for the future load as well. TTE to take this into account for new design.
03/09/17 – TTE sent new email to County with recommended Chiller sizes. We are awaiting a new submittal from Trane to confirm physical sizes of new chiller shells. TTE awaiting County’s best guess of combined chiller load during summer. Emcor says they do not have this information on file.
03/15/17 – Based on conversation with HamCo and ThermalTech load calcs we are gong to proceed with a 350 ton chiller as the base chiller for this project. This gives us enough future machine to handle the addition of the 9th and 10th floors if and when those spaces are renovated.
03/21/17 – LCC equation complete. TTC writing bid specs for vendors.
04/26/16 – First set of quotes have arrived at ThermalTech. We are beginning the LCC process.
05/24/17 – Met with County presented LCC results for this project. Narrowed selection down to two chillers to be used in bid. ThermalTech conducted more fieldwork for the replacement project.
06/12/17 – THP – Provided a proposal for structural engineering design services to Thermal Tech.
06/26/17 – We are down to two chillers one for the base bid with standard refrigerant and one with future refrigerant. We will design around these two machines for the bid. We have finally reached three workable designs on the cooling towers, one from Marley, one from BAC and one from Evapco. The Marley solution consists of four cooling towers, Evapco and BAC consists of 2 towers. We are trying for a July bid opening for this project or maybe early August.
06/28/17 – County okay with bidding out two chiller options and three cooling tower options, ThermalTech to bid out in July and start construction ASAP knowing that CT may be done in 2018.
07/12/17 – THP notified via email proposal is acceptable.
07/24/17 – Chilled Water pumps are scheduled to arrive next week. TEI to work with DK to get installed in a fashion that works for this job. Change Order submittal must be discussed. Three cooling tower options have been design, one base bid, two alternates. After field work from last week, even these designs are very tight. TEI to meet with THP and discuss. A new review set is ready for County review.
07/27/17 – Penthouse roof under warranty with Tremco and its the white aggregate. THP will start design on the dock and roof next week.
08/15/17 – THP provided review letter for proposed new cooling tower loading.
08/18/17 – THP provided review drawings for structural and waterproofing repairs at the loading dock.
08/30/17 – Project is out for bid. Prebid scheduled for 1:30pm on 9/14/17. THP issued structural and waterproofing loading dock repair documents.
09/18/17 – THP issued addendum 01 questions responses to TEI.
09/25/17 – Addendums 1,2,3 have been issued on this project. We have extended the Bid Opening to Oct 4, 2017.
10/24/17 – Bid were opened and both bids (Peck and DK) were over budget. A post bid analysis was completed by ThermalTech and delivered to Facilities on 10/24 for their use. Several items were identified as being out of scope, under estimated, or omitted from the original scope. The analysis found that the project probably is about $380,000 to $420,000 above the original scope estimate.
11/28/17 – ThermalTech collected new bids on alternate equipment in an effort to get this project in under the original budget. We have received pricing from Trane, JCI and McQuay. We are working on an LCC cost to go along with this equipment to see hat impact the different equipment types have on the project.
11/30/17 – Final prices due by 12/1 at 2pm to Bert. TTC and Barth to take care of.
01/25/18 – ThermalTech to rebid the chillers out as air cooled chillers for LCC then bid this out as a new project.
02/21/18 – ThermalTech preparing new bid for LCC Air Cooled Chillers. We should have this out for bid in a week. TEI has fully developed the PID for the new air cooled chiller scenario and reviewed internally to make sure it works being a departure from the existing system and it will work. Some piping modifications will take place, a bypass will be moved and the chilled water return on the air handlers on the first floor must be reconnected in a different place. TEI also laid out new chillers on the roof structure and they are too big to fit into the old penthouse area so we developed two alternate scenarios where they can sit on the roof between the existing DX rooftop units. TEI contacted THP for SOSS to redesign the system for the new locations. TEI will submit revised plans and cost to Facilities within a week.

by TEI TEI

CAB – Auditor’s 3rd Floor Renovation

07/25/16 – Facilities contacts GBBN and IPM about Kick-off meeting on this project.
08/17/16 – Kick-off meeting established for 08/24/16.
08/24/16 – Kick-off Meeting held at CAB Third Floor. Walk through of space held after meeting conclusion. Facilities forwards two carpet replacement options for separate but simultaneous carpet replacement project. GBBN contacts TEI for MEP scope input and options for the SOSS authoring.
08/24/16 – GBBN forwards documents to ThermalTech for investigation into the MEP scope and SOSS involvement.
08/25/16 – IPM to perform field verification of existing system.
08/29/16 – IPM performs additional field verification of the east side of the Third Floor.
08/30/16 – GBBN receives three re-carpeting Options for the various areas of the Third Floor from Facilities.
08/31/16 – GBBN authors and delivers the SOSS to Facilities.
09/01/16 – IPM transmits first preliminary layout of the Auditor’s Office Renovation per the owner’s directives. Auditor responds the layout still needs more work.
09/06/16 – IPM transmits second preliminary layout of the Auditor’s Office Renovation per the owner’s directives. Auditor responds the layout still needs more work.
09/07/16 – IPM transmits third preliminary layout of the Auditor’s Office Renovation per the owner’s directives. Auditor responds the layout still needs more work.
09/13/16 – IPM transmits fourth preliminary layout of the Auditor’s Office Renovation per the owner’s directives. Auditor responds the layout still needs more work.
09/21/16 – Auditor’s Office meeting with IPM, GBBN and Facilities held to discuss final approved layout and logistics for implementing project with various vendors.
10/12/16 – GBBN and Facilities hold discussions on renumbering and renaming the various spaces affected by this renovation.
10/14/16 – GBBN completes construction documents and forwards to ThermalTech to begin MEP design.
10/10/16 – Facilities emails GBBN about conversations with Paula Drake with the Auditor Office. She would like to move forward with the project without the expedited walk-through process. The new proposed gypsum board partitions will not be added until next year. Spectra Flooring is directed to remove the 5th floor VCT flooring from their quote.
10/24/16 – IPM and GBBN are contacted that recent retirements in the Auditor Office will require additional changes to the layout. IPM contacts Auditor for meeting to discuss.
10/25/16 – IPM completes latest revisions to the preliminary layout and obtains approval from Auditor. GBBN updates ThermalTech with new drawing and we will work with Architect to get EC dwg prepared. HVAC dwg will be given to GBBN for inclusion on their dwgs if possible.
12/21/16 – ThermalTech releases permitting drawings for this renovation.
12/29/16 – Renovations are submitted for permitting and work is underway.
01/25/17 – Permit was received today for this project.
01/26/17 – Paula Drake emails to indicate consolation of more tax maps were accomplished and therefore one short map cabinets could be dismantled.
01/27/17 – Planes Moving indicates Spectra and Planes will mobilize on Saturday for installation of carpet and furniture moving.
01/28/17 – IPM reports the moving and carpet installers arrived late. Carpet installers indicated they were not aware demolition of existing carpeting was involved.
01/29/17 – IPM indicated carpet was installed but base was not completed.
02/16/17 – Facilities calls GBBN to inform that Paula Drake has requested several Owner Changes. These changes are extensive enough to require an Engineering Change to the City of Cincinnati. GBBN to inquire with IPM and TEI for estimated fees for making the drawing and as-built changes, including reimbursable expenses for Engineering Change.
03/15/17 – GBBN coordinates scope of services for Engineering Changes with ThermalTech and IPM. This includes application and plan review fees for IBI and City of Cincinnati Buildings and Inspections. The SOSS is authored and submitted to Facilities.
05/25/17 – Project on hold pending new direction.
07/31/17 – City of Cincinnati Building Inspector called GBBN to discuss status of Permit Inspections. GBBN indicated Owner has stopped work based on potential changes that would require Additional Services for an Engineering Change. Inspector indicated he would indicate work was on hold.
10/18/17 – Interior Project Management indicates Auditor has requested some additional work to be performed on next phase.
02/22/18 – Project ready to archive. City has approved permit and CO. Extend permit within 10 days prior to expiration.